### v1

# Axi-majoron for almost everything

Abstract
The details of the minimal cosmological standard model (MCSM) proposed in Ref. [1] are discussed. The model is based on the scale-symmetry and the global Peccei-Quinn(PQ) symmetry with a key assumption that the latter is broken only in the gravity sector in a scale-invariant manner. We show that the model provides a quite simple unified framework for the unknown history of the universe from inflation to the epoch of big-bang nucleosynthesis, simultaneously addressing key puzzles of high energy theory and cosmology: (i) the origin of scales, (ii) primordial inflation, (iii) matter-antimatter asymmetry, (iv) tiny neutrino masses, (v) dark matter, and (vi) the strong CP-problem. Scale symmetry can be exact, and the Planck scale is dynamically generated. The presence of Gauss-Bonnet term may safely retain dangerous non-perturbative symmetry-breaking effects negligible, allowing large-field trans-Planckian inflation along the PQ-field. Iso-curvature perturbations of axi-majorons are suppressed. A sizable amount of PQ-number asymmetry is generated at the end of inflation, and conserved afterwards. Domain wall problem is absent due to the PQ-number asymmetry. Baryogenesis can be realized by either inflationary Affleck-Dine mechanism or spontaneous leptogenesis thanks to the PQ-number asymmetry, or by resonant leptogenesis. Dark matter can be purely cold axi-majorons from the mis-alignment contribution only with the symmetry-breaking scale of $\mathcal{O}(10^{12}){\rm GeV}$. Hot axi-majorons from the decay of the inflaton become a natural source for a sizable amount of dark radiation. Inflationary gravitational waves are expected to have information about some masse parameters of the left-handed and the right-handed neutrinos, thanks to the presence of an early matter-domination era driven by at least a long-lived right-handed neutrino species.
1 Introduction
In modern physics, the subatomic world is described by the standard model(SM) of particle physics [2, 3], and the large scale associated with our observable universe is governed by Einstein gravity. They are associated with two vastly different scales, the electroweak scale for the former and the Planck scale for the latter, as set by various experiments. These scales appear as dimensionful constants in the corresponding theories. The presence of such constants may mean the existence of a defining higher level theory in which dimensionful parameters are absent [^1] Scale-invariance has been discussed in various contexts including cosmological applications (see [4] for example and references therein).. Then, all scales should be generated dynamically, and differences among scales would be controlled by dimensionless parameters. Question still remains though for the origin of the dimensionless parameters. They might require extra-structure of spacetime including our 4-$D$ spacetime.
Recently, there have been several works discussing dynamical generation of scales in a scale invariant version of gravity and extensions of the SM [5, 6, 7, 8]. An interesting finding is that the Planck scale can be dynamically generated even without a Coleman-Weinberg(CW)-type symmetry-breaking [9], thanks to the conservation of scale current which has a kernel settling down to a constant [5, 8]. Low energy scales, such as the electroweak scale, can be generated in connection with the field responsible for the generation of the Planck scale, or separately via CW-mechanism in the presence of matter fields other than SM fields. Scale-invariance can be maintained even at quantum level, provided that the renormalization scale (which usually appears as an explicit dimensionful parameter) is replaced by a fundamental scalar field. Preservation of scale-invariance decouples the massless dilaton from the other fields, hence makes the theory free from fifth-force constraints [10, 11]. This implies that a fully scale-invariant theory can be consistent with low energy phenomenology.
Meanwhile, any theory trying to explain the current state of the universe should be able to accommodate a phase of primordial inflation  [^2] It was found that in the scenario of an emergent gravity with scale-invariance a slow-roll inflation can arise quite naturally [7, 12]. However, connection with low energy physics was missing., generation of matter-antimatter asymmetry, proper candidate(s) of dark matter, neutrino masses inferred by neutrino flavor oscillation data [13], and probably a solution to the strong CP-problem such as invisible axion [14, 15, 16, 17]. All of these big puzzles of cosmology and phenomenology can not be solved within the SM only, and definitely demand a theory beyond the standard model(BSM). In this regard, the idea of unifying axion and majoron [18] is quite interesting, since it can be realized in a scale-invariant manner, simultaneously address the tiny neutrino masses, and provide an axion solution to the strong CP-problem in a very simple structure. Then, it is tempting to see if the idea can be combined with inflation and generation of matter-antimatter asymmetry while keeping the scale-invariance.
In Ref. [1], we proposed a minimal scale-invariant $U(1)_{\rm PQ}$-symmetric extension of the SM [19, 20], dubbed as ‘the minimal cosmological standard model(MCSM)’, to address the aforementioned issues altogether other than the notorious Cosmological Constant(CC) problem. In this work, we show the details of the MCSM. It is a DFSZ-type axion model [16, 17] unified with the singlet majoron model, in which Peccei-Quinn(PQ) field not only provides the axion solution, but it is also responsible for the generation of the masses of right-handed neutrinos. Hence, we call the associated pseudo-Goldstone boson as ‘axi-majoron’. In the sense of unifying axion and majoron, our scenario is similar to the earlier works such as Refs. [18] and [21, 22], based on KSVZ [14, 15] and DFSZ [16, 17] axion models, respectively. Also, in regard of providing a unified framework addressing aforementioned puzzles, one can find works such as Refs. [23, 24]. Our scenario is different from those works in various aspects including dynamical generation of the Planck scale, inflationary physics, generation of matter-antimatter asymmetry and physics of axions. In particular, assuming $U(1)_{\rm PQ}$-symmetry to be broken in the gravity sector only [^3] There are some earlier works such as Refs. [25] and [26] where the PQ symmetry-breaking non-minimal gravitational interaction was considered. In Ref. [25], the authors payed attention only to the impact of such an interaction on the axion-isocurvature perturbations which can be suppressed by a temporary axion mass much larger than the expansion rate during inflation thanks to the interaction. In Ref. [26], angular motion was removed by setting initial position of the field at the valley of the angular potential for simplicity. In both works, the PQ-field are located at the angular minimum position during inflation. In our case, the angular motion and the non-zero angular displacement of the PQ-field at the end of inflation play the critical role for the post inflation cosmology., we show that the dynamics of the inflaton along axi-majoron direction can generate a large amount of PQ-number asymmetry which is eventually converted to the baryon number asymmetry via the seesaw sector. It is a kind of Affleck-Dine(AD) mechanism [27] realized by Peccei-Quinn field. Interestingly, such an asymmetry corresponding to a motion of axi-majoron allows not only AD-mechanism but also spontaneous leptogenesis. Even in the case of PQ-number asymmetry not being enough, our scenario allows resonant leptogenesis [28] to work. Also, very intriguingly, in regard to the axion solution to the strong CP-problem, the symmetry-breaking terms in the gravity sector are harmless at least at perturbative level, and may be so even at non-perturbative level thanks to topological curvature-terms such as Gauss-Bonnet [29]. Iso-curvature perturbations are suppressed, and the domain-wall problem is absent due to the angular motion of PQ-field. Also, a sizable amount of dark radiation from hot axi-majorons can be obtained naturally, and may ameliorate the Hubble tension [30] [31]. The standard thermal bath is recovered by decays of at least a long-lived non-relativistic right-handed neutrino species. Because of the presence of an early matter domination era, primordial gravitational waves from inflation are expected to have a characteristic spectral feature, containing information of the masses of the lightest left-handed and the lightest right-handed neutrinos. In combinations with dark radiation and inflationary observables, it may provide a unique way to probe such physical quantities. Our scenario provides thus a very minimal, unique unified framework for the unknown history of the universe from inflation to the BBN epoch.
This paper is organized as follows. In Section 2, the model is described. In Section 3, inflation with regard to dynamics of the homogeneous mode, evolutions of perturbations, and reheating are discussed. In Section 4, we show how leptogenesis works in our scenario via various mechanisms thanks to the dynamics of PQ-field. In Section 5, dark matter and dark radiation from axi-majorons are briefly discussed. In Section 6, conclusions are drawn.
2 The Model
We assume that scale-invariance is a fundamental symmetry and preserved in both of the matter and the gravity sectors, and the Planck scale is dynamically generated by spontaneous breaking of scale symmetry. We also assume that the global Peccei-Quinn symmetry, $U(1)_{\rm PQ}$, is preserved but only in the matter sector, since gravity does not seem to respect global symmetries. Based on this symmetry, in order to realize the axion solution to the strong CP-problem and the seesaw mechanism for neutrino masses and mixings in a minimal unified setup, we consider a scale-invariant Type-II two Higgs doublet extension of the SM [32] [^4] KSVZ-type realization such as Ref. [23] is also possible, and heavy extra quarks can contribute to dark matter relic density in this case. with the seesaw sector in the form of the singlet majoron model [33]. The assignment of PQ-charges to fields is shown in Table 1.
Charge Field $Q_{L}$ $u_{R}$ $d_{R}$ $\ell_{L}$ $e_{R}$ $\nu_{R}$ $H_{1}$ $H_{2}$ $\Phi$
$q_{\rm PQ}$ 3/2 1/2 1/2 3/2 1/2 1/2 1 -1 1
Table 1: Charge assignment of fields under $U(1)_{\rm PQ}$ and $U(1)_{B-L}$ symmetries. $\nu_{R}$ is the right-handed neutrino(RHN) field, $H_{1}=\left(H_{1}^{+},\ H_{1}^{0}\right)^{T}$ and $H_{2}=\left(H_{2}^{+},\ H_{2}^{0}\right)^{T}$ are down- and up-type Higgs doublets, respectively. $\Phi$ is the Peccei-Quinn field. Flavor indices of matter fields were suppressed.
There are four scalar fields as the minimal set: a real scalar $\chi$ which is responsible for generating Planck scale via spontaneous breaking of scale symmetry, Type-II two Higgs doublets, and one complex scalar which plays the role of the Peccei-Quinn field, denoted here as $\Phi$.
Our model is minimal in the sense of the number of fields introduced to accommodate various phenomenological/cosmological/theoretical necessities. It might be compared to the model discussed in Ref. [23] for example. The difference is how the axion solution is realized, either DFSZ-type or KSVZ-type. While our scenario introduces only one extra Higgs doublet as a minimal DFSZ axion model, the compared model introduces two extra heavy quarks as a minimal KSVZ axion model. Esthetically, we think our model is much nicer in regard of its symmetrical structure and the very attractive efficiency achieved in such a simple structure with a minimal set of fields.
Instead of considering all these fields explicitly, we deliver our argument with a simplified version, having only one SM Higgs doublet $H$. The presence of the second Higgs doublet which is essential for a realization of DFSZ axion scenario does not affect the main argument. Paying attention to only these fields and their relevant interactions, we characterize our model by the following action555 The largrangian of the gravity sector can have terms involving $H_{1}H_{2}$ such as $\tilde{R}\left(H_{1}^{\dagger}H_{2}\pm{\rm h.c.}\right)$. Also, in the matter sector, there can be terms such as $\Phi^{2}\left(H_{1}^{\dagger}H_{2}\right)+{\rm h.c.}$. We ignore these terms in this work since we will be interested in the case in which the couplings of $H_{1}$ and $H_{2}$ to $\Phi$ are negligible and hence those Higgs fields are likely to stay at the symmetric phase before the electroweak phase transition (EWPT). We will discuss else where the case where initially $H_{1}$ and $H_{2}$ have very large field values during inflation. :
$S_{\rm G}$ $=$ $-\frac{1}{2}\int d^{4}x\sqrt{-{\tilde{g}}}{\tilde{R}}\left[\xi_{\chi}\chi^{2}+2\xi_{h}|H|^{2}+2\xi_{\phi}|\Phi|^{2}\right.$ (2.1)
$\left.+\xi_{+}\left(\Phi^{2}+{\rm c.c.}\right)-i\xi_{-}\left(\Phi^{2}-{\rm c.c.}\right)\right]\ ,$
$S_{\rm M}$ $=$ $\int d^{4}x\sqrt{-{\tilde{g}}}\left[{\tilde{g}}^{\mu\nu}\left(\frac{1}{2}\partial_{\mu}\chi\partial_{\nu}\chi+\partial_{\mu}H^{\dagger}\partial_{\nu}H+\partial_{\mu}\Phi^{\dagger}\partial_{\nu}\Phi\right)-{\tilde{V}}\right]\ ,$ (2.2)
where all $\xi$s are assumed to be real positive definite in order for the effective $M_{\rm P}^{2}$ to be positive in the large field region along any direction, and the potential $\tilde{V}$ is given by
${\tilde{V}}$ $=$ ${\tilde{V}}_{S}+{\tilde{V}}_{F}\ ,$ (2.3)
${\tilde{V}}_{S}$ $=$ $\frac{\lambda_{\chi}}{4}\chi^{4}+\lambda_{h}|H|^{4}+\lambda_{\phi}|\Phi|^{4}$ (2.4)
$-\frac{1}{2}\lambda_{\chi h}\chi^{2}|H|^{2}-\frac{1}{2}\lambda_{\chi\phi}\chi^{2}|\Phi|^{2}-\lambda_{h\phi}|H|^{2}|\Phi|^{2}\ ,$
${\tilde{V}}_{F}$ $=$ $\frac{1}{2}y_{N}\Phi^{*}\overline{\nu_{R}^{c}}\nu_{R}+y_{\nu}\overline{\ell}_{L}{\tilde{H}}\nu_{R}+{\rm h.c.}\ .$ (2.5)
where ${\tilde{H}}\equiv i\tau_{2}H^{*}$ with $\tau_{2}$ being the second Pauli matrix [^6] In the full model, ${\tilde{H}}$ should be replaced to $\tilde{H_{2}}$, the conjugate of the up-type Higgs doublet., and the flavor indices were suppressed. All $\lambda$s in Eq. (2.4) are assumed to be positive real free parameters. Note that, thanks to the Yukawa interaction $y_{N}$, the spontaneous breaking of $U(1)_{\rm PQ}$ is responsible for Majorana masses of RHN fields. For this reason, we will call the associated pseudo-Goldstone boson as axi-majoron.
[…]
Decomposing $H$ and $\Phi$ as
$H^{T}(x)=(0,h(x)/\sqrt{2}),\ \ \ \Phi(x)=\phi(x)e^{i\theta(x)}/\sqrt{2}=\left(\phi_{r}(x)+i\phi_{i}(x)\right)/\sqrt{2},$ (2.6)
and introducing a new mass scale $M_{\rm P}$ to be fixed later, the action can be expressed as
$S_{\rm G}$ $=$ $-\frac{1}{2}\int d^{4}x\sqrt{-{\tilde{g}}}{\tilde{R}}M_{\rm P}^{2}\left[\frac{\xi_{\chi}\chi^{2}}{M_{\rm P}^{2}}+\frac{\xi_{h}h^{2}}{M_{\rm P}^{2}}+\frac{\xi_{\phi}\phi^{2}}{M_{\rm P}^{2}}+\frac{\xi_{a}\phi^{2}\cos(2\theta)}{M_{\rm P}^{2}}\right]\ ,$ (2.7)
$S_{\rm M}$ $=$ $\int d^{4}x\sqrt{-{\tilde{g}}}\left[\frac{1}{2}{\tilde{g}}^{\mu\nu}\left(\partial_{\mu}\chi\partial_{\nu}\chi+\partial_{\mu}h\partial_{\nu}h+\sum_{a}\partial_{\mu}\phi^{a}\partial_{\nu}\phi^{a}\right)-{\tilde{V}}\right]\ ,$ (2.8)
where in Eq. (2.7) we redefined $\Phi$ as
$\Phi\to e^{i\theta_{0}}\Phi$ (2.9)
so that
$\left(\xi_{+}-i\xi_{-}\right)e^{i2\theta_{0}}=\sqrt{\xi_{+}^{2}+\xi_{-}^{2}}\equiv\xi_{a}$ (2.10)
and $\phi^{a}=\left(\phi_{r},\phi_{i}\right)^{T}$ in Eq. (2.8). We define a frame-function $\Omega$ as
$\Omega$$\equiv$ $\frac{\xi_{\chi}\chi^{2}}{M_{\rm P}^{2}}+\frac{\xi_{h}h^{2}}{M_{\rm P}^{2}}+\frac{\xi_{\phi}\phi^{2}}{M_{\rm P}^{2}}\left(1+\alpha\cos(2\theta)\right)$ (2.11)
$=$ $\frac{\xi_{\chi}\chi^{2}}{M_{\rm P}^{2}}+\frac{\xi_{h}h^{2}}{M_{\rm P}^{2}}+\frac{\xi_{r}\phi_{r}^{2}}{M_{\rm P}^{2}}+\frac{\xi_{i}\phi_{i}^{2}}{M_{\rm P}^{2}}\ ,$
where $\alpha\equiv\xi_{a}/\xi_{\phi}$, $\xi_{r}\equiv\xi_{\phi}+\xi_{a}$, $\xi_{i}\equiv\xi_{\phi}-\xi_{a}$ ($\xi_{r}\geq\xi_{i}$). We assume
$\xi_{r}>\xi_{i}\quad\Leftrightarrow\quad\xi_{\phi}>\xi_{a}\quad\Leftrightarrow\quad\alpha<1$ (2.12)
in order for $\Omega$ to be positive definite even in the large field region, $\xi_{\phi}^{1/2}\phi>M_{\rm P}$, which will be of our interest for a primordial inflation. Without loss of generality, the scalar potential ${\tilde{V}}_{S}$ can be expressed as  [^7] The presence of the second Higgs doublet makes the shape of the potential complicated, but such details are not important in the discussion of the present work.
${\tilde{V}}_{S}(\chi,h,\phi)$ $=$ $\frac{\lambda_{\chi}^{\prime}}{4}\chi^{4}+\frac{\lambda_{\phi}^{\prime}}{4}\left(\phi^{2}-\zeta_{\phi\chi}\chi^{2}\right)^{2}+\frac{\lambda_{h}}{4}\left(h^{2}+\zeta_{h\phi}\phi^{2}-\zeta_{h\chi}\chi^{2}\right)^{2}\ .$ (2.13)
Hence, once the vacuum expectation value (VEV) of $\chi$ is fixed so that $\xi_{\chi}^{1/2}\langle\chi\rangle=M_{\rm P}\simeq 2.44\times 10^{18}{\rm GeV}$(see Ref. [8] for example for dynamical generation of Planck scale in a scale invariant theory), the VEVs of $\phi$ and $h$ at low energy can be fixed to
$\phi_{0}$ $=$ $\sqrt{\zeta_{\phi\chi}\xi_{\chi}^{-1}}M_{\rm P}=\mathcal{O}(10^{9-12})\ {\rm GeV}\ ,$ (2.14)
$h_{0}$ $=$ $\sqrt{\zeta_{h\chi}\xi_{\chi}^{-1}M_{\rm P}^{2}-\zeta_{h\phi}\phi_{0}^{2}}\simeq 246\ {\rm GeV}\ ,$ (2.15)
and ${\tilde{V}}_{S}$ has the familiar form,
${\tilde{V}}_{S}(h,\phi)=V_{0}+\frac{\lambda_{h}}{4}\left(h^{2}-h_{0}^{2}\right)^{2}+\frac{\lambda_{\phi}^{\prime}}{4}\left(\phi^{2}-\phi_{0}^{2}\right)^{2}\ .$ (2.16)
Note that, if $\lambda_{h\phi}=0$(or equivalently $\zeta_{h\phi}=0$) which we assume in this work [^8] Actually, it is enough to have $\lambda_{h\phi}\ll\left(\lambda_{h\chi}/\xi_{\chi}\right)\times\left(M_{\rm P}/\phi_{0}\right)^{2}$ which is not a special condition., the VEVs are determined by the couplings to $\chi$ only, i.e., $\chi$ is responsible for all scales.
In the theory characterized by the action in Eqs. (2.1) and (2.2), Planck scale is determined by the sum of the VEVs of all scalar fields. The sum, however, should be dominated by the VEV of $\chi$, since the other fields are supposed to have much smaller VEVs at low energy. The determination of Planck scale implies breaking of scale-invariance. Since there are no explicit symmetry-breaking terms, this should be achieved by spontaneous symmetry-breaking. In Ref. [8], it was shown that Planck scale can emerge naturally in scale-invariant theories even without CW-type dimensional transmutation. This is due to the fact that the conserved scale-current has a kernel which eventually settles down to a constant after some general expansion in a very early epoch, at least well before the last about $60$ inflationary $e$-folds. Scale-anomaly can compromise this picture, but it was shown that, once the renormalization scale, which usually appears as a dimensionful parameter, is replaced by a field, scale symmetry becomes exact, and the existence of a constant kernel remains valid [34]. The constant value of the kernel defines an ellipse in the multi-dimensional scalar-field space 999 Quantum effects on $\xi_{i}$ can change the elliptic configuration [8], but the impact on our discussion is minor. Hence, we ignore the effects in this work., and eventually determines the Planck scale at low energy, once all fields settle down to their true vacuum position.
Applying the analysis of Ref. [8] to our scenario, when $h\lll\chi,\phi_{r},\phi_{i}$ in the very early epoch including inflation era, we are left with a three-dimensional field space. In terms of new field variables defined as  [^10] The index $\alpha$ should not be confused with $\alpha$ defined in Eq. (2.11).
$X_{\alpha}\equiv\xi_{\alpha}\varphi_{\alpha}^{2}\ ,$ (2.17)
with $\alpha=(1,2,3)$, $\xi_{\alpha}=\left(\xi_{\chi},\xi_{r},\xi_{i}\right)$, and $\varphi_{\alpha}=\left(\chi,\phi_{r},\phi_{i}\right)$, the constant kernel is given by
$K=\frac{1}{2}\sum_{\alpha}\kappa_{\alpha}\varphi_{\alpha}^{2}\ ,$ (2.18)
where $\kappa_{\alpha}\equiv 1+6\xi_{\alpha}$. If the initial condition and the dynamics of fields are such that
$\kappa_{\chi}\chi^{2}\gg\kappa_{r}\phi_{r}^{2}+\kappa_{i}\phi_{i}^{2}$ (2.19)
during the whole evolution of fields, $K$ would always be dominated by the contribution of $\chi$. In terms of $X_{\alpha}$, the condition Eq. (2.19) corresponds to
$X_{1}\gg\frac{\xi_{1}}{\kappa_{1}}\left[\left(\frac{\kappa_{2}}{\xi_{2}}\right)X_{2}+\left(\frac{\kappa_{3}}{\xi_{3}}\right)X_{3}\right]\ .$ (2.20)
For $\xi_{a}\ll\xi_{\phi}$(i.e., $\alpha\ll 1$) which will be of our interest, $\xi_{2}\approx\xi_{3}$ and we can take $\xi_{\alpha}$s such as
$\frac{\xi_{1}}{\kappa_{1}}\lll\frac{\xi_{2}}{\kappa_{2}}\approx\frac{\xi_{3}}{\kappa_{3}}\ \left(<\frac{1}{6}\right)\ ,$ (2.21)
while Eq. (2.20) is always satisfied. Then, one can have
$X_{1}(0)\ll X_{2}(0)+X_{3}(0)$ (2.22)
as the initial condition for the primordial inflation responsible for our universe. If ${\tilde{V}}_{S}$ is dominated by the $\lambda_{\phi}\phi^{4}$-term during inflation, in the slow-roll limit of the primordial inflation, the evolution equation in the Jordan frame is approximated as  [^11] Actually, since the dynamics is constrained to be on the ellipse defined in Eq. (2.18), one of the equations in Eq. (2.26) is redundant. [8]
$\left(\begin{array}[]{c}X_{1}^{\prime}\\ X_{2}^{\prime}\\ X_{3}^{\prime}\end{array}\right)$ $=$ $8\left(\frac{\sum_{i}X_{i}}{\sum_{i}\kappa_{i}X_{i}}\right)\frac{1}{{\tilde{V}}_{S}}$ (2.26)
$\times$ $\left(\begin{array}[]{cc}\xi_{1}X_{1}\left[\kappa_{2}\left({\tilde{V}}_{22}+{\tilde{V}}_{23}\right)+\kappa_{3}\left({\tilde{V}}_{23}+{\tilde{V}}_{33}\right)\right]\\ -\xi_{2}\left[\kappa_{1}X_{1}\left({\tilde{V}}_{22}+\frac{{\tilde{V}}_{23}}{2}\right)-\kappa_{3}\left(X_{2}\left(\frac{{\tilde{V}}_{23}}{2}+{\tilde{V}}_{33}\right)-X_{3}\left({\tilde{V}}_{22}+\frac{{\tilde{V}}_{23}}{2}\right)\right)\right]\\ -\xi_{3}\left[\kappa_{1}X_{1}\left({\tilde{V}}_{33}+\frac{{\tilde{V}}_{23}}{2}\right)+\kappa_{2}\left(X_{2}\left(\frac{{\tilde{V}}_{23}}{2}+{\tilde{V}}_{33}\right)-X_{3}\left({\tilde{V}}_{22}+\frac{{\tilde{V}}_{23}}{2}\right)\right)\right]\end{array}\right)\ ,$ (2.30)
where $\prime$” denotes the derivative with respect to $N=\ln a(t)$, and
${\tilde{V}}_{22}\equiv\frac{\lambda_{\phi}}{4}\phi_{r}^{4},\ {\tilde{V}}_{23}\equiv\frac{\lambda_{\phi}}{2}\phi_{r}^{2}\phi_{i}^{2},\ {\tilde{V}}_{33}\equiv\frac{\lambda_{\phi}}{4}\phi_{i}^{4}\ .$ (2.31)
Note that we have ${\tilde{V}}_{ij}/{\tilde{V}}\lesssim 1$ for $\phi_{r}\gtrsim\phi_{i}$. Hence, from Eq. (2.26), it is clear that, for $\xi_{1}\lll 1/8$ with the initial condition satisfying Eqs. (2.20) and (2.22), $X_{1}$ barely changes during the dynamics of $X_{2,3}$ moving toward minima along each direction. In such a circumstance, if $X_{3}=0$ as a simple possibility, the EOM of $X_{2}$ can be trivially integrated as
$X_{2}(N)-X_{2}(0)+\frac{\kappa_{1}-\kappa_{2}}{\kappa_{2}}X_{1}(0)\ln\left[\frac{X_{1}(0)+X_{2}(N)}{X_{1}(0)+X_{2}(0)}\right]=-8\xi_{2}\left(\frac{\kappa_{1}}{\kappa_{2}}\right)X_{1}(0)N\ .$ (2.32)
It is valid for $X_{2}(N)\gtrsim X_{1}(0)$ with the change of $X_{1}$ ignored, and gives $e$-folds as
$N\approx\frac{\kappa_{2}}{8\xi_{2}\kappa_{1}}\frac{X_{2}(0)-X_{2}(N)}{X_{1}(0)}\ \xrightarrow[]{X_{2}(0)\gg X_{2}(N)}\ \frac{\kappa_{2}}{8\xi_{2}\kappa_{1}}\frac{X_{2}(0)}{X_{1}(0)}>\frac{3}{4}\frac{X_{2}(0)}{X_{1}(0)}\ .$ (2.33)
Hence, one can obtain sufficiently large amount of $e$-folds if $X_{2}(0)\gg X_{1}(0)$. Note that, since $X_{1}$ barely changes for a sufficiently small $\xi_{1}$ while $X_{2,3}$ becomes negligible in the end, it is possible to assume an approximation of a fixed Planck scale whose value given by
$M_{\rm P}\simeq\sqrt{\xi_{\chi}}\chi_{0}\simeq\sqrt{\frac{2\xi_{\chi}K}{1+6\xi_{\chi}}}\ .$ (2.34)
Then, the physics becomes the same as that of non-minimal gravitational couplings of scalar fields other than $\chi$ with the fixed Planck scale. We take this approximation in the subsequent discussion to describe the dynamics $X_{2,3}$ in the more transparent and clear, conventional way.
Once the Planck scale is (nearly) fixed in this way, we can follow the conventional way of converting the Jordan frame to the Einstein frame. The frame-function in Eq. (2.11) with the Higgs field ignored is now expressed as
$\Omega\equiv 1+\frac{\xi_{\phi}\phi^{2}}{M_{\rm P}^{2}}\left(1+\alpha\cos(2\theta)\right)=1+\frac{\xi_{r}\phi_{r}^{2}}{M_{\rm P}^{2}}+\frac{\xi_{i}\phi_{i}^{2}}{M_{\rm P}^{2}}\ .$ (2.35)
Under a Weyl rescaling,
$g_{\mu\nu}=\Omega{\tilde{g}}_{\mu\nu}$ (2.36)
with $g_{\mu\nu}$ the metric in the Einstein frame, the action of the gravity part obtains a form in the Einstein gravity [35],
$S_{\rm G}=-\frac{1}{2}\int d^{4}x\sqrt{-g}M_{\rm P}^{2}\left[R-\frac{3}{2}g^{\mu\nu}\left(\partial_{\mu}\ln\Omega\right)\left(\partial_{\nu}\ln\Omega\right)\right]\ ,$ (2.37)
while the action of matter sector is given by
$S_{\rm M}=\int d^{4}x\sqrt{-g}\left\{\frac{1}{2\Omega}\left[\left(\partial h\right)^{2}+\sum_{a}\left(\partial\phi^{a}\right)^{2}\right]-\frac{{\tilde{V}}}{\Omega^{2}}\right\}\ .$ (2.38)
The action for the scalar fields part can now be written as
$S_{\rm scalar}=\int d^{4}x\sqrt{-g}\left[\frac{1}{2}\mathcal{K}_{ab}g^{\mu\nu}\partial_{\mu}\phi^{a}\partial_{\nu}\phi^{b}-U\right]\ ,$ (2.39)
where the symmetric field-space metric $\mathcal{K}_{ab}$ is given by
$\mathcal{K}_{ab}=\Omega^{-2}\left[\Omega\delta_{ab}+\frac{3M_{\rm P}^{2}}{2}\frac{\partial\Omega}{\partial\phi^{a}}\frac{\partial\Omega}{\partial\phi^{b}}\right]\equiv\Omega^{-2}K_{ab}$ (2.40)
and $U\equiv{\tilde{V}}_{S}(h,\phi)/\Omega^{2}$. At low energy with $h,\phi\lll M_{\rm P}$, $\Omega\to 1$ and the kinetic terms return to the canonical form. The shape of the potential $U$ is shown in Fig. 1.
Figure 1: Potential energy density $U$ with $\xi_{\phi}=1$, $\lambda_{\phi}=6\times 10^{-10}$, and $\alpha=0.5$ as a function of $\phi$ for various $\theta$ (left), and as a function of $\theta$ for various $\phi$ (right).
The non-minimal symmetry-breaking gravitational interaction, $\xi_{a}$-term contributes to the mass of the axi-majoron, $a_{\phi}\equiv\theta\phi$, the angular mode of $\Phi$ field for a given non-zero $\phi$. The low-energy definition of axi-majoron differs due to the symmetry-breaking non-zero VEVs of the PQ-charged two Higgses, but such a change does not affect our argument. In the Einstein frame, in the vicinity of true vacuum position, the mass-squared can be read as
$m_{a}^{2}\equiv\frac{\partial^{2}U}{\phi^{2}\partial\theta^{2}}=8\alpha\xi_{\phi}\frac{U}{M_{\rm P}^{2}}\ .$ (2.41)
Hence, depending on $\alpha\ \xi_{\phi}$, it can be comparable to or smaller than the expansion rate. Note that, even though it is from a symmetry-breaking term, this contribution to the axi-majoron mass does not disturb the axion solution, since it becomes negligible at low energy, but helps suppress the iso-curvature perturbations and play the key role for baryogenesis after inflation, as discussed in the subsequent sections.
3 Inflation along $\Phi$ direction
We consider an initial condition, $h\ll M_{\rm p}\lesssim\phi$. Consistency with our assumption discussed in the previous section, i.e.,
$\kappa_{r}\phi_{r,\rm ini}^{2}+\kappa_{i}\phi_{i,\rm ini}^{2}\ll\kappa_{\chi}\chi_{\rm ini}^{2}$ (3.1)
with $M_{\rm P}\simeq\sqrt{\xi_{\chi}}\chi_{\rm ini}$ in this analysis constrains the initial position of $\phi$ as
$\sqrt{\xi_{\phi}}\phi_{\rm ini}\ll\sqrt{\frac{1+6\xi_{\chi}}{\xi_{\chi}}\times\frac{\xi_{\phi}}{1+6\xi_{\phi}\left(1+\alpha\cos\left(2\theta\right)\right)}}M_{\rm P}\ \xrightarrow[]{\xi_{\chi}\lll 1\ \&\ \xi_{\phi}\gg 1}\ \frac{M_{\rm P}}{\sqrt{6\xi_{\chi}}}\ .$ (3.2)
Therefore, for a sufficiently small $\xi_{\chi}$ with $\xi_{\phi}\gtrsim\mathcal{O}(1)$, it is possible to have $\sqrt{\xi_{\phi}}\phi_{\rm ini}\gg M_{\rm P}$ while Planck scale is nearly fixed, and we assume this is the case. In this work, we will be interested in $\xi_{\phi}=\mathcal{O}(1-10^{2})$, concerning various cosmological aspects of this model. Hence, the unitarity-cutoff may be slightly lower than the Planck scale, but it is well above any energy and mass scales involved in the field dynamics. Also, even though the field excursion during inflation spans over the Planck scale as will be discussed subsequently, it may present no problem as long as scale-invariance is preserved.
In the large field region around or over the Planck scale, the kinetic terms of the scalar fields $\phi_{r}$ and $\phi_{i}$ in Einstein frame are neither canonical nor diagonal. In order to analyze the dynamics of fields in such a curved field-space, we resort to the conventional method for multi-field inflation in a curved field-space [36, 37]. From now on we work in the basis of $\phi^{a=1,2}\equiv(\phi_{r},\phi_{i})^{T}$ unless otherwise stated. We define the Levi-Civita connection, covariant field-derivative $\nabla_{a}$, and covariant time-derive in the field space as
$\Gamma_{ab}^{c}$ $=$ $\frac{1}{2}\mathcal{K}^{cd}\left(\mathcal{K}_{da,b}+\mathcal{K}_{db,a}-\mathcal{K}_{ab,d}\right)\ ,$ (3.3)
$\nabla_{a}\phi^{c}$ $=$ $\partial_{a}\phi^{c}+\Gamma_{ab}^{c}\phi^{b}\ ,$ (3.4)
$\nabla_{t}$$\equiv$ $\dot{\phi}^{a}\nabla_{a}\ ,$ (3.5)
where $\partial_{a}\equiv\partial/\partial\phi^{a}$, $\mathcal{K}^{ab}$ is the inverse of $\mathcal{K}_{ab}$, and $\mathcal{K}_{ab,c}\equiv\partial\mathcal{K}_{ab}/\partial\phi^{c}$. Each component of the metric can be found in a straightforward manner and given in Appendix A in terms of $K_{ab}$ for convenience. The speed of the inflaton is given by $\dot{\phi}_{I}\equiv\sqrt{\mathcal{K}_{ab}\dot{\phi}^{a}\dot{\phi}^{b}}$. Then, the vectors tangent and normal to the trajectory of the field dynamics are respectively given by
$T^{a}$ $=$ $\frac{\dot{\phi}^{a}}{\dot{\phi}_{I}}$ (3.6)
$N_{a}$ $=$ $\sqrt{\mathcal{K}}\epsilon_{ab}T^{b}$ (3.7)
where $\mathcal{K}$ is the determinant of the metric $\mathcal{K}_{ab}$, and $\epsilon_{ab}$ is the antisymmetric Levi-Civita tensor.
3.1 Dynamics of the homogeneous mode
For the flat FLRW metric, $g_{\mu\nu}={\rm diag}\left(1,-a^{2}(t),-a^{2}(t),-a^{2}(t)\right)$, the equations of motions of the background fields are given by
$0=\nabla_{t}\dot{\phi}^{a}+3H\dot{\phi}^{a}+\mathcal{K}^{ab}U_{a}\ ,$ (3.8)
where $U_{a}\equiv\partial U/\partial\phi^{a}$, and Friedman equation is given by
$3H^{2}M_{\rm P}^{2}=\frac{1}{2}\mathcal{K}_{ab}\dot{\phi}^{a}\dot{\phi}^{b}+U\ .$ (3.9)
Contracting the equation of motion of the field vector $\phi^{a}$ in Eq. (3.8) with either $T_{a}$ or $N_{a}$, one obtains the equations of motions along the inflaton trajectory and orthogonal to that as
00 0 $=$ $\ddot{\phi}_{I}+3H\dot{\phi}_{I}+U_{T}$ (3.10)
$\nabla_{t}T^{a}$ $=$ $-\frac{U_{N}N^{a}}{\dot{\phi}_{I}}\equiv-\dot{\theta}_{T}N^{a}\ ,$ (3.11)
where $U_{T}\equiv T^{a}U_{a}$, $U_{N}\equiv N^{a}U_{a}$, and $\dot{\theta}_{T}\equiv U_{N}/\dot{\phi}_{I}$ is the turning rate of the inflaton trajectory.
One can solve numerically these equations of motion of fields straightforwardly. However, it is instructive to see a limiting case which can be proven simpler to understand. For that, one can take a look at the case of $\alpha\lll 1$ for whose(?) approximate analytic expressions can be obtained as follows. The frame function in $\left(\phi,\theta\right)$-basis is approximated as
$\Omega\approx 1+\frac{\xi_{\phi}\phi^{2}}{M_{\rm P}^{2}}\ .$ (3.12)
The field-space metric is approximated as
$K_{11}$$\approx$ $1+\frac{\xi_{\phi}\left(1+6\xi_{\phi}\right)\phi^{2}}{M_{\rm P}^{2}}\ ,$
$K_{22}$$\approx$ $1+\frac{\xi_{\phi}\left[1+6\xi_{\phi}\left(\alpha\cos(2\theta)\right)^{2}\right]\phi^{2}}{M_{\rm P}^{2}}\ ,$
$K_{12}$$\approx$ $6\left(\frac{\xi_{\phi}\phi}{M_{\rm P}}\right)^{2}\alpha\cos\left(2\theta\right)\ll 1\ .$ (3.13)
In order to have a clear picture, let us ignore the mixing term $K_{12}$ for the moment [^12] This term, however, is included in numerical analysis.. Then, the canonically normalized fields can be defined as
$d\varphi_{1}$$\equiv$ $\Omega^{-1}\sqrt{1+\frac{\xi_{\phi}\left(1+6\xi_{\phi}\right)\phi^{2}}{M_{\rm P}^{2}}}d\phi\ ,$ (3.14)
$d\varphi_{2}$$\equiv$ $\Omega^{-1/2}\phi d\theta\ .$ (3.15)
Note that, as expected, the field $\varphi_{1}$ plays the same role as the canonically normalized inflaton of the SM Higgs inflation [38]. In the large field region, $\xi_{\phi}^{1/2}\phi\gg M_{\rm P}$, we have
$\phi$ $\simeq$ $\frac{M_{\rm P}}{\sqrt{\xi_{\phi}}}\exp\left[\sqrt{\frac{\xi_{\phi}}{1+6\xi_{\phi}}}\frac{\varphi_{1}}{M_{\rm P}}\right]\ ,$ (3.16)
$\theta$ $\simeq$ $\frac{\varphi_{2}}{M_{\rm P}/\sqrt{\xi_{\phi}}}\ ,$ (3.17)
where $\varphi_{2}$ should be regarded as a periodic field, i.e.,
$\varphi_{2}=\varphi_{2}+2\pi\left(\frac{M_{\rm P}}{\sqrt{\xi_{\phi}}}\right)\ .$ (3.18)
In terms of these approximate canonical fields, the potential is given by
$U(\varphi_{1},\varphi_{2})$ $=$ $\frac{\lambda_{\Phi}M_{\rm P}^{4}}{4\xi_{\phi}^{2}}\left[1+\alpha\cos\left(\frac{2\sqrt{\xi_{\phi}}\varphi_{2}}{M_{\rm P}}\right)+\exp\left[-2\sqrt{\frac{\xi_{\phi}}{1+6\xi_{\phi}}}\frac{\varphi_{1}}{M_{\rm P}}\right]\right]^{-2}\ .$ (3.19)
[…]
$H_{I}\simeq\frac{\lambda_{\Phi}^{1/2}M_{\rm P}}{2\sqrt{3}\xi_{\phi}}$ (3.20)
in the large field region.
Meanwhile, from the derivatives along $\varphi_{2}$ direction, one finds (see Appendix B)
$\frac{M_{\rm P}}{U}\frac{\partial U}{\partial\varphi_{2}}$ $=$ $4\sqrt{\xi_{\phi}}\alpha\sin\left(\frac{2\sqrt{\xi_{\phi}}\varphi_{2}}{M_{\rm P}}\right)\ ,$ (3.21)
$\frac{M_{\rm P}^{2}}{U}\frac{\partial^{2}U}{\partial\varphi_{2}^{2}}$ $=$ $8\xi_{\phi}\alpha\cos\left(\frac{2\sqrt{\xi_{\phi}}\varphi_{2}}{M_{\rm P}}\right)\ ,$ (3.22)
$\frac{M_{\rm P}^{2}}{U}\frac{\partial^{2}U}{\partial\varphi_{1}\partial\varphi_{2}}$ $=$ $\frac{\frac{12\xi_{\phi}}{\sqrt{1+6\xi_{\phi}}}\alpha\exp\left[2\sqrt{\frac{\xi_{\phi}}{1+6\xi_{\phi}}}\frac{\varphi_{1}}{M_{\rm P}}\right]\sin\left(\frac{2\sqrt{\xi_{\phi}}\varphi_{2}}{M_{\rm P}}\right)}{\left[1+\left(1+\alpha\cos\left(\frac{2\sqrt{\xi_{\phi}}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[2\sqrt{\frac{\xi_{\phi}}{1+6\xi_{\phi}}}\frac{\varphi_{1}}{M_{\rm P}}\right]\right]^{2}}\ .$ (3.23)
From Eq. (3.22), the mass-squard along $\varphi_{2}$-direction is found to be
$m_{\varphi_{2}}^{2}\approx 24\alpha\xi_{\phi}H_{I}^{2}\cos\left(\frac{2\sqrt{\xi_{\phi}}\varphi_{2}}{M_{\rm P}}\right)$ (3.24)
Hence, if $\alpha\sim\alpha_{\rm c}$ with
$\alpha_{\rm c}\equiv 1/(24\xi_{\phi}),$ (3.25)
the mass along $\varphi_{2}$ direction (i.e., angular direction in the complex field space of $\Phi$) is comparable to the expansion rate given in Eq. (3.20). This implies that, in the large field region where the potential along the radial direction (i.e., along $\varphi_{1}$ or $\phi$) becomes quite flat, if $\alpha\gtrsim\alpha_{\rm c}$, the dynamics of field would be initially along $\varphi_{2}$ toward the vicinity of $\varphi_{2}\sim 0$ and afterward moves nearly along $\varphi_{1}$ toward the origin, as will be shown by numerical analysis. On the other hand, if $\alpha\lll\alpha_{\rm c}$, the trajectory will be nearly radial towards the origin. That is, the trajectory of the field would look like one of the following three cases in $(\phi_{r},\phi_{i})$ space:
  • C-I : $\sqrt{|m_{\varphi_{2}}^{2}|}\gtrsim H_{I}$ ($\Leftrightarrow\alpha\gtrsim\alpha_{\rm c}$) - $\varphi_{2}\to 0$ (angular motion) and then $\varphi_{1}\to 0$ (radial motion).
  • C-II : $|m_{\varphi_{2}}^{2}|\sim|m_{\varphi_{1}}^{2}|$ ($\Leftrightarrow\alpha\ll\alpha_{\rm c}$) - a curvilinear trajectory (mixture of angular and radial motion).
  • C-III : $|m_{\varphi_{2}}^{2}|\ll|m_{\varphi_{1}}^{2}|$ ($\Leftrightarrow\alpha\lll\alpha_{\rm c}$) - a nearly radial trajectory.
In Fig. 2, trajectories of field dynamics are shown for different values of $\alpha$ with a fixed initial position as an example.
Figure 2: Trajectories of the field dynamics for $\alpha=10^{-4},10^{-3},10^{-2},10^{-1}$ from left to right with $\xi_{\phi}=1$, $\lambda_{\phi}=6\times 10^{-10}$, $\phi_{\rm ini}=20M_{\rm P}$, and $\theta_{\rm ini}=\pi/16,\pi/4,7\pi/16$(from left to right).
Note that the angular motion in the case of $\alpha=0.1$ is closer to the origin relative to the case of $\alpha=10^{-2}$. This is because, as $\alpha$ becomes sufficiently large, the effect of the mixing between $\varphi_{1}$ and $\varphi_{2}$ becomes significant, if $\theta_{\rm ini}=\mathcal{O}(1)$ as inferred in Eq. (3.23)
As it can be seen from the dynamics of the inflaton with $\theta_{\rm ini}\neq 0$, a salient and interesting feature of our hybrid axi-majoron inflation is that the PQ field $\Phi$ can have angular motion even after inflation, if $\alpha$ is neither too large nor too small. Fig. 3 shows the time-dependence of $\theta$ and its time-derivative.
Figure 3: The time dependence of the phase $\theta$ and its time-derivative for $\alpha=10^{-3}$ with $\xi_{\phi}=1$, $\lambda_{\phi}=6\times 10^{-10}$, $\phi_{\rm ini}=20M_{\rm P}$, and $\theta_{\rm ini}=\pi/4$. $H_{\rm ref}\equiv\sqrt{U(\phi_{\rm ini},\theta_{\rm ini})}/\sqrt{3}M_{\rm P}\approx H_{I}$. The vertical dotted line indicates the end of inflation.
The angular momentum in a co-moving volume is conserved after inflation, since the symmetry-breaking effect becomes negligible as $\phi$ becomes smaller than thePlanck scale. This corresponds to the asymmetry of $U(1)_{\rm PQ}$-charged particle $\Phi$ as in the case of Affleck-Dine mechanism [27]  [^13] In Ref. [39], generation of Peccei-Quinn number asymmetry and subsequent baryogenesis was discussed, but in a different realization.. The asymmetry can be transferred to the visible sector, as the baryon number asymmetry, via the usual way of Affleck-Dine mechanism and/or the mechanism of spontaneous baryogenesis [40]. More details will be discussed in Section 4.
3.2 Perturbations
For the estimations of perturbations [^14] There are perturbations of $\chi$ field too. However, for $\xi_{\chi}\ll\xi_{r,i}$ the background dynamics is nearly along $\Phi$ with barely changing $\chi$. In addition, in a simplified two field case with negligible kinetic mixing which is the case of $\xi_{\chi}\lll 1$, one can see that the growing rate of iso-curvature perturbation is negative (i.e., suppression instead of growth) in the slow-roll and slow-turn limit [41]. Hence, we can ignore the perturbations of $\chi$ in our discussion. , we take the Newtonian gauge in an isotropic universe for the metric perturbations, i.e., the perturbed metric is given by
$ds^{2}=\left(1+2\psi\right)dt^{2}-a(t)^{2}\left(1-2\psi\right)\delta_{ij}dx^{i}dx^{j}\ .$ (3.26)
In terms of the Mukhanov-Sasaki variable for gauge invariant perturbations [42, 43],
$Q^{a}\equiv\delta\phi^{a}+\frac{\dot{\phi}^{a}}{H}\psi$ (3.27)
with $Q_{T}\equiv T_{a}Q^{a}$ and $Q_{N}\equiv N_{a}Q^{a}$, adiabatic and isocurvature perturbations can be defined as
$\mathcal{R}$$\equiv$ $\frac{H}{\dot{\phi}_{I}}Q_{T}\ ,$ (3.28)
$\mathcal{S}$$\equiv$ $\frac{H}{\dot{\phi}_{I}}Q_{N}\ .$ (3.29)
Also, one can define slow-roll functions as [44, 36]
$\epsilon$$\equiv$ $-\frac{\dot{H}}{H^{2}}=\frac{\dot{\phi}_{I}^{2}}{2H^{2}}$ (3.30)
$\eta^{a}$$\equiv$ $-\frac{\nabla_{t}\dot{\phi}^{a}}{H\dot{\phi}_{I}}=\eta_{T}T^{a}+\eta_{N}N^{a}$ (3.31)
with
$\eta_{T}$$\equiv$ $T_{a}\eta^{a}=-\frac{\ddot{\phi}_{I}}{H\dot{\phi}_{I}}\ ,$ (3.32)
$\eta_{N}$$\equiv$ $N_{a}\eta^{a}=\frac{U_{N}}{H\dot{\phi}_{I}}=\frac{\dot{\theta}_{T}}{H}\ .$ (3.33)
They are related to the conventional slow-roll parameters defined in terms of the potential in single field scenarios as
$\epsilon$ $\simeq$ $\epsilon_{U_{T}}$ (3.34)
$\eta_{T}$ $\simeq$ $\eta_{U_{T}}-\epsilon_{U_{T}}$ (3.35)
where $\epsilon_{U_{T}}\equiv\left(M_{\rm P}U^{\prime}/U\right)^{2}/2$ and $\eta_{U_{T}}\equiv M_{\rm P}^{2}U^{\prime\prime}/U$ with the derivatives taken with respect to the inflaton field.
The quadratic order action for the perturbations is given by [45]
$S_{2}$ $=$ $\frac{1}{2}\int d^{4}x\sqrt{-g}\left[\frac{\dot{\phi}_{I}^{2}}{H^{2}}g^{\mu\nu}\left(\partial_{\mu}\mathcal{R}\right)\left(\partial_{\nu}\mathcal{R}\right)+g^{\mu\nu}\left(\partial_{\mu}Q_{N}\right)\left(\partial_{\nu}Q_{N}\right)\right.$ (3.36)
$\left.-m_{\rm eff}^{2}Q_{N}^{2}+4\dot{\phi}_{I}\eta_{N}\dot{\mathcal{R}}Q_{N}\right]\ .$
It gives the EOMs of the perturbations as
¨$\ddot{\mathcal{R}}+\left(3+2\epsilon-2\eta_{T}\right)H\mathcal{R}+\frac{k^{2}}{a^{2}}\mathcal{R}$ $=$ $-2\dot{\theta}_{T}\frac{H}{\dot{\phi}_{I}}\left[\dot{Q}_{N}+\left(3-\eta_{T}-2\epsilon+\frac{\dot{\eta}_{N}}{H\eta_{N}}\right)HQ_{N}\right]$ (3.37)
$\ddot{Q}_{N}+3H\dot{Q}_{N}+\frac{k^{2}}{a^{2}}Q_{N}+m_{\rm eff}^{2}Q_{N}$ $=$ $2\dot{\theta}_{T}\frac{\dot{\phi}_{I}}{H}\dot{\mathcal{R}}$ (3.38)
where $k$ is the comoving wave number associated with cosmological scales. The effective mass of the fluctuations orthogonal to the inflaton direction is given by
$m_{\rm eff}^{2}=m^{2}-\dot{\theta}_{T}^{2}\ ,$ (3.39)
where
$m^{2}=U_{NN}+\epsilon H^{2}\mathbb{R}$ (3.40)
with $U_{NN}\equiv N^{a}N^{b}\nabla_{a}\nabla_{b}U$ and $\mathbb{R}$ being the Ricci scalar associated with the field-space metric $\mathcal{K}_{ab}$.
For numerical analysis, we took $\xi_{\phi}=1$, $\alpha=10^{-3}$, $\lambda_{\phi}=6\times 10^{-10}$, $\phi_{\rm ini}=20M_{\rm P}$, and $\theta_{\rm ini}=\pi/4$ as an illustrative set of parameters.
Figure 4: Left: The turning rate, $\dot{\theta}_{T}(t)$, for $t\leq t_{e}$. Right: The effective mass-squared $m_{\rm eff}^{2}$ associated with iso-curvature perturbations for the same parameter set as the one used in Fig. 3. The vertical dotted line corresponds to $52$ $e$-folds to the end of inflation at $t=t_{e}\simeq 352H_{\rm ref}^{-1}$. Roughly, it corresponds to the time when the co-moving scale associated with the pivot scale of the Planck mission exits the horizon during inflation.
In Fig. 4, $\dot{\theta}_{T}$ and $m_{\rm eff}^{2}$ are depicted. From the left panel of the figure, one can see that, $|\dot{\theta}_{T}|=\mathcal{O}(10^{-3})\times H_{I}$, i.e., the turning rate of the inflaton trajectory is quite small. Also, the right panel shows that $m_{\rm eff}^{2}$ is not negligible and positive at least for the last about 60 $e$-folds of inflation, leading to a suppression of iso-curvature perturbations. As a result, the source terms in both of Eqs. (3.37) and (3.38) (i.e., the right-hand sides of the equations) are negligible, and there is no sizable impact of a two-dimensional curvilinear trajectory of the inflaton on the evolution of adiabatic and isocurvature perturbations associated with cosmological scales of the present universe. Even though the inflaton trajectory involves axi-majoron direction which has anomalous couplings to non-Abelian gauge fields of the SM, the perturbations of gauge fields are not enhanced because the velocity of axi-majoron during inflation is maintained to be smaller than the expansion rate by at least a couple of orders of magnitude during the whole relevant $e$-folds, as shown in Fig. 3. Non-Gaussianities are not enhanced [46] either. In Fig. 5, we depict adiabatic (left) and iso-curvature perturbations (right) for comoving scales exiting horizon at $N\approx 37$(dashed red) and $52$(solid blue) as examples.
Figure 5: Adiabatic and iso-curvature perturbations of the mode $k_{N}=a_{N}H_{I}$ for $N=37$(dashed red) and $52$(solid blue) with the same parameter set as the one used in Fig. 3.
In a multidimensional field-space, if the direction of the motion of the inflaton is changed during inflation, the sound speed of perturbations can be affected due to the mixing of kinetic terms [47]. Specifically, the sound speed $c_{s}$ is given by [48]
$c_{s}^{-2}(k)=1+\frac{4\dot{\theta}_{T}^{2}}{\left(k/a\right)^{2}+m_{\rm eff}^{2}}\ .$ (3.41)
However, from Fig. 4 as an illustrative example and subsequent discussion, we see that the change of $c_{s}$ is negligible, and set $c_{s}=1$. Then, the amplitude of the power spectrum, the spectral index, and the tensor-to-scalar ratio are given by
$P_{R}$ $=$ $\frac{1}{8\pi^{2}\epsilon}\times\frac{H^{2}}{M_{\rm P}^{2}}\ ,$ (3.42)
$n_{s}$ $=$ $1-4\epsilon+2\eta_{T}\ ,$ (3.43)
$r_{T}$ $=$ $16\epsilon\ .$ (3.44)
For the potential given in Eq. (3.19) with $\alpha\ll 1$, the relations between slow-roll parameters and $e$-folds $N_{e}$ in the SM Higgs-inflation [38] are expected still to hold at least approximately. The latter is given by [38]
$N_{e}(\phi)\approx\frac{3}{4}\frac{\phi^{2}-\phi_{e}^{2}}{M_{\rm P}^{2}/\xi_{\phi}}$ (3.45)
as the $e$-folds from the end of inflation at $\phi_{e}\approx\left(4/3\right)^{1/4}M_{\rm P}/\sqrt{\xi_{\phi}}$. Slow-roll parameters are then given by
$\epsilon$ $\simeq$ $\frac{3}{4}N_{e}^{-2}\ ,$ (3.46)
$\eta_{T}$ $\simeq$ $N_{e}^{-1}\ ,$ (3.47)
leading to
$P_{R}$ $\simeq$ $\frac{\lambda_{\phi}N_{e}^{2}}{72\pi^{2}\xi_{\phi}^{2}}\ ,$ (3.48)
$n_{s}$ $\simeq$ $1-2N_{e}^{-1}\ ,$ (3.49)
$r_{T}$ $\simeq$ $12N_{e}^{-2}\ .$ (3.50)
Density perturbations of the universe extracted from various observations including CMB observations lead to $P_{R}\simeq 2.1(\pm 0.03)\times 10^{-9}\ (1\sigma\ {\rm CL})$, $n_{s}=0.9658\pm 0.004\ (1\sigma\ {\rm CL})$, and $r_{T}<0.068\ (2\sigma\ {\rm CL})$ for the pivot scale, $k_{*}=0.05/{\rm Mpc}$ [49]. Hence, Eq. (3.48) constrains $\lambda_{\phi}$ and $\xi_{\phi}$ to satisfy a relation,
$\lambda_{\phi}\sim 6\times 10^{-10}\xi_{\phi}^{2}\left(\frac{52}{N_{e}}\right)^{2}\ .$ (3.51)
where we used $N_{e}=52$ as a representative value associated with $k_{*}$. In Table 2, we show a result of numerical estimations for some sets of parameters set as
$\lambda_{\phi}\equiv\lambda_{\phi}^{\rm ref}\times\xi_{\phi}^{2},\ \alpha=\frac{10^{-3}}{\xi_{\phi}},\ \phi_{\rm ini}=\frac{20M_{\rm P}}{\sqrt{\xi_{\phi}}},\ \theta_{\rm ini}=\frac{\pi}{4}\ ,$ (3.52)
with $\lambda_{\phi}^{\rm ref}\equiv 6\times 10^{-10}$ and several different values of $\xi_{\phi}$ but $N_{e}=52$ fixed.
$\xi_{\phi}$ $P_{R}\times 10^{9}$ $n_{s}$ $r_{T}\times 10^{3}$
1 2.067 0.9622 4.738
10 2.387 0.9620 4.128
$10^{2}$ 2.424 0.9619 4.066
$10^{3}$ 2.428 0.9619 4.060
Table 2: Examples of numerical estimations for inflationary observables with parameters set as Eq. (3.52).
The result is in a good agreement with approximations in Eq. (3.48) – (3.50), and can match well with observations by adjusting $\lambda_{\phi}$. The sizable difference between the cases of $\xi_{\phi}=1$ and the other much larger values may be understood from the dependence of $U$ on $\sqrt{\xi_{\phi}/(1+6\xi_{\phi})}$ in Eq. (3.19). For convenience, in the subsequent argument we use $\lambda_{\phi}$ parameterized as in Eq. (3.52).
In our scenario, there are five parameters affecting the observables associated with inflation in the complex field space of PQ-field:
$\xi_{\phi},\lambda_{\phi},\alpha,\theta_{\rm ini},\phi_{\rm ini}\ .$ (3.53)
Once $\xi_{\phi}$ as a free parameter is chosen such that $\xi_{\phi}\gtrsim\mathcal{O}(1)$ [^15] As long as $\lambda_{\phi}\lll 1$ with $\xi_{\phi}$ given by the relation Eq. (3.51), the expansion rate during inflation and the maximal energy scale right after inflation are always well below the scale of the unitarity cutoff, $\Lambda_{\rm cut}\sim M_{\rm P}/\sqrt{\xi_{\phi}}$ [50, 51, 52]. , $\lambda_{\phi}$ is nearly fixed as long as $\alpha$ is sufficiently smaller than unity. Dependence of observables on the other parameters for given $\xi_{\phi}$ and $\lambda_{\phi}$ is shown in Fig. 6.
Figure 6: $P_{R}$(left) and $n_{s}$(right) as functions of $\theta_{\rm ini}$ for various values of $\alpha$ with $\xi_{\phi}=1$, $\lambda_{\phi}=6\times 10^{-10}$, $\phi_{\rm ini}/M_{\rm P}=10({\rm top}),20({\rm bottom})$.
It turns out that inflationary observables are barely affected by $\theta_{\rm ini}$ for $\sqrt{\xi_{\phi}}\phi_{\rm ini}\gtrsim 10M_{\rm P}$ as long as $\alpha\xi_{\phi}\lesssim\mathcal{O}(10^{-2})$ except the case of $\theta_{\rm ini}$ close to $\pi/2$ with $\sqrt{\xi_{\phi}}\phi_{\rm ini}\sim 10M_{\rm P}$ when $\alpha\xi_{\phi}\sim 10^{-2}$. Note that, if angular dynamics is not strong, for a given $e$-folds associated with the pivot scale of observations, the spectral index is nearly fixed irrespective of initial position of the field dynamics. The amplitude of power spectrum has only minor dependence on the initial position, and the dependence can be compensated by adjusting $\lambda_{\phi}$ without affecting the spectral index. This is because, while the former is a function of $e$-folds only, the latter is a function of $e$-folds and energy density of inflation. On the other hand, if $\alpha\xi_{\phi}\gtrsim\mathcal{O}(10^{-1})$, the observables varies a lot, depending on $\theta_{\rm ini}$, but $\phi_{\rm ini}$-dependence becomes weak if it is sufficiently larger than the minimal value to accommodate about 60 $e$-folds.
3.3 Reheating after inflation
Once inflation ends, after a transient period, the energy density of $\phi^{4}$ model evolves in the same way as radiation [53]. In the left panel of Fig. 7, the evolution of the scale factor and expansion rate are shown.
Figure 7: Left: Evolutions of the scale factor (dashed blue) and $\sqrt{H_{e}/H(t)}$ (solid red) as functions of cosmic time $t$ from the end of inflation for the same parameter set as the one used in Fig. 3. The solid and dashed lines are nearly overlapped, implying $H(t)\propto a(t)^{-2}$. The dot-dashed green line is of $\left(t/t_{e}\right)^{1/2}$. The dotted green line is the same as the dot-dashed green line but just up-shifted for a clear comparison to red or dashed blue line. Right: Evolution of the energy density of $\Phi$ (solid red). The dashed green line corresponds to $t^{-2}$-behavior shown for a comparison.
One can see that $H(t)\propto a(t)^{-2}$ from the end of inflation. However, note that these lines do not match to the dot-dashed green line (lower diagonal line) which corresponds to $\left(t/t_{e}\right)^{1/2}$. Their time dependence becomes equal only some time after inflation. The actual value of the scale factor (dashed blue) is larger than the value of the lower green line by about one order of magnitude. The energy density of $\Phi$ scales like radiation not right after inflation but some time later after getting through a short suppression period (see the right panel of the figure).
For $\phi$ well below the Planck scale, the scalar potential is approximated as $U\simeq{\tilde{V}}_{S}$. It can be expressed as a familiar form given in Eq. (2.16). For simplicity, we consider the case of $\lambda_{\phi}^{\prime}=\lambda_{\phi}$ (i.e., $\zeta_{h\phi}=0\ \Leftrightarrow\lambda_{h\phi}=0$). Then, the mass of $\phi$-quanta for a given background value is given by
$m_{\phi}=\sqrt{3\lambda_{\phi}}\phi\ \xrightarrow[]{\phi\to\phi_{0}}\ \sqrt{2\lambda_{\phi}}\phi_{0}\ .$ (3.54)
For $\lambda_{h\phi}=0$ the production of the SM Higgs from either preheating processes or decays of $\phi$ particles is negligible. For sufficiently small $y_{N}$ which will be discussed in the subsequent discussion, the production of RHNs via preheating is also not efficient [54]. Hence, initially there can be population of $\phi$-particles and axi-majorons ($a_{\phi}$) only  [^16] There are thermal particles with a temperature $T\sim H_{\rm e}/(2\pi)$ at the end of inflation [55]. We ignore this contribution which does not affect our argument. . As shown in Ref. [56], the self-preheating effect on $\phi$-particles is inefficient. The population of $a_{\phi}$ is only from the perturbative decay of $\phi$-particles through the kinetic term of $a_{\phi}$. Moreover, if $\Phi$ has non-zero angular momentum (i.e., a phase rotation), the preheating effect will be suppressed since $\phi$ does not pass through the origin during its oscillations and the violation of adiabaticity which is necessary for an efficient preheating becomes weak as a result. However, there might be a possibility of efficient energy transfer via axi-majoron’s anomalous couplings to non-Abelian gauge fields due to large velocity of axi-majoron after inflation [57]. Since the velocity of axi-majoron is periodically changing as shown in Fig. 3, in regard of the polarization state of gauge field which is amplified the situation is similar to the case of parametric resonance in a conformal scalar theory [56]. The details will be investigated in other work. In any case, as long as the eventual decay of $\phi$-particle is the dominant source of the energy of the universe and the interaction of $\phi$ to RHNs is small enough to avoid the possibility of thermal-trapping, our argument is barely affected. Hence, in this work we ignore the effect of preheating.
In our scenario, $U(1)_{\rm PQ}$ symmetry is broken only in the gravity sector, and its effect is suppressed by $\alpha\xi_{\phi}\phi^{2}/M_{\rm P}^{2}$. As a result, $a_{\phi}$ is lighter than $\phi$ for $\phi\lesssim M_{\rm P}$ with $\alpha\xi_{\phi}\ll 1$, and $\phi$ can always decay to $a_{\phi}$s via the kinetic term of $a_{\phi}$. The partial decay rate of the channel is given by
$\Gamma_{a_{\phi}}^{\phi}=\frac{m_{\phi}^{3}}{32\pi\phi^{2}}\approx\left\{\begin{array}[]{ll}\left(3\sqrt{3}/32\pi\right)\lambda_{\phi}^{3/2}\phi&:\phi>\phi_{\times}\\ \left(2\sqrt{2}/32\pi\right)\lambda_{\phi}^{3/2}\phi_{0}&:\phi<\phi_{\times}\end{array}\right.\ ,$ (3.55)
where $\phi$ and $\phi_{\times}(\equiv\sqrt{2}\phi_{0})$ are respectively the oscillation amplitude and the point where the potential energy of $\phi$-field is the same as the one at the origin. The inflaton should be able to decay to RHNs too in order to recover the standard thermal background consisting of mainly the SM particles. For the mass of a RHN mass-eigenstate $N_{i}$, $m_{N_{i}}=y_{N_{i}}\phi_{0}/\sqrt{2}$, the channel is open only if
$y_{N_{i}}<\sqrt{\lambda_{\phi}}\ .$ (3.56)
From the seesaw formula for the mass parameters of the left-handed neutrinos [58, 59]  [^17] At low energy below the electroweak scale, $y_{\nu}$ should be replaced to the one obtained from a bi-unitarity transformation associated with physical mass-eigenstates. We ignore such details here, since it does not make changes of the main point of our argument. Hence, we assumed $y_{\nu}$ is flavor-diagonal for simplicity.,
$m_{\nu_{i}}\approx\frac{y_{\nu_{i}}^{2}v_{h}^{2}}{2m_{N_{i}}}\ .$ (3.57)
Then the constraint in Eq. (3.56) is translated as a constraint on $y_{\nu_{i}}$, which can be expressed as
$y_{\nu_{i}}^{2}$$\approx$ $2\sqrt{b_{i}\lambda_{\phi}}\left(\frac{m_{\nu_{i}}\phi_{0}}{v_{h}^{2}}\right)\ ,$ (3.58)
where we have used $b_{i}\equiv y_{N_{i}}^{2}/2\lambda_{\phi}$ for later convenience. For simplicity, we assume that $m_{N_{i}}$s and $m_{\nu_{i}}$s have the same hierarchy. Note that neutrino oscillation data is consistent with the massless lightest left-handed neutrino, which means that in principle $m_{\nu_{1}}$(correspondingly $y_{\nu_{1}}$) associated with the lightest mass eigenstate can be arbitrarily small. The decay rate $\phi$ to a RHN mass-eigenstate $N_{i}$ is given by
$\Gamma_{N_{i}}^{\phi}=b_{i}\left(1-2b_{i}\right)^{3/2}\Gamma_{a_{\phi}}^{\phi}\ .$ (3.59)
where we used $m_{\phi}=\sqrt{2\lambda_{\phi}}\phi$ for $\phi\gtrsim\phi_{0}$. Only these channels are relevant for the decay of $\phi$. Hence, the total decay rate of $\phi$ is given by
$\Gamma^{\phi}=\Gamma_{a_{\phi}}^{\phi}+\sum_{i}\Gamma_{N_{i}}^{\phi}\ ,$ (3.60)
and the branching ratio of $N_{i}$-channel is given by
${\rm B}_{i}=\frac{b_{i}}{1+\sum_{j}b_{j}}<\frac{1}{3}\ ,$ (3.61)
where the phase space factor of the RHN-channel was ignored.
For $\phi\gtrsim\phi_{\times}$ the universe is driven by a radiation-like energy component. In this case, the decay can happen effectively once $H\approx\Gamma^{\phi}/2$ is satisfied. If the subdominant angular motion is ignored, the expansion rate after inflation can be expressed such as
$H=\frac{\lambda_{\phi}^{1/2}\phi^{2}}{2\sqrt{3}M_{\rm P}}\propto a^{-2}\ ,$ (3.62)
where $\phi\propto a^{-1}$ with $a$ being the scale factor of the expansion of the universe was used for the scaling behavior of $H$. Then, taking $\Gamma^{\phi}\approx\Gamma_{a_{\phi}}^{\phi}$, one finds that the decay of $\phi$ takes place when $\phi=\phi_{*}$ with
$\frac{\phi_{*}}{\phi_{\times}}\approx\frac{9\lambda_{\phi}}{32\sqrt{2}\pi}\frac{M_{\rm P}}{\phi_{0}}\simeq\left(\frac{\xi_{\phi}}{10^{2}}\right)^{2}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)\ .$ (3.63)
which is valid only for $\phi_{*}>\phi_{\times}$.
If $\phi_{*}<\phi_{\times}$, decays start only after the field falls into the vicinity of the low energy true vacuum position, i.e., $\phi_{0}$. In such a case, the field amplitude at the epoch of decays is found as follows. With a redefined field ${\tilde{\phi}}\equiv\phi-\phi_{0}<\phi_{0}$, the potential of $\phi$ can be approximated as
$V(\phi)\approx\lambda_{\phi}\left[\phi_{0}^{2}{\tilde{\phi}}^{2}+\phi_{0}{\tilde{\phi}}^{3}+\frac{1}{4}{\tilde{\phi}}^{4}\right]\ \xrightarrow[]{{\tilde{\phi}}\ll\phi_{0}}\ \frac{1}{2}m_{\phi}^{2}{\tilde{\phi}}^{2}\ .$ (3.64)
The energy density of $\phi$ is then given by
$\rho_{\phi}\approx\frac{1}{2}m_{\phi}^{2}{\tilde{\phi}}^{2}\ ,$ (3.65)
where ${\tilde{\phi}}$ is understood as the amplitude of oscillation, and $m_{\phi}=\sqrt{2\lambda_{\phi}}\phi_{0}$. In this regime, the universe is in a matter-domination era driven by non-relativistic massive $\phi$-particles [^18] There is no kination era driven by the kinetic anergy caused by angular motion of $\Phi$. This is because the phase velocity of $\Phi$, i.e., $\dot{\theta}$ scales as $a^{-1}$ for $\phi\gtrsim\phi_{\times}$ and $a^{-3}$ for $\phi\lesssim\phi_{\times}$ in our scenario. Its energy contribution is always sub-dominant relative to the contribution from the radial mode.. Hence, requiring $H=(2/3)\Gamma^{\phi}$, one finds the oscillation amplitude ${\tilde{\phi}}_{*}$ at the epoch of the decay of $\phi$ as
${\tilde{\phi}}_{*}\approx\frac{\sqrt{6}\lambda_{\phi}}{24\pi}M_{\rm P}\simeq 4.7\times 10^{7}{\rm GeV}\times\xi_{\phi}^{2}\ ,$ (3.66)
and the corresponding expansion rate is given by
$H_{*}\approx\frac{\sqrt{2}\lambda_{\phi}^{3/2}}{24\pi}\phi_{0}\simeq 3.5\times 10^{-4}{\rm GeV}\times\xi_{\phi}^{3}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.67)
One may worry about the possibility of holding $\phi$ around the origin due to thermal effect before $\phi$ falls below $\phi_{\times}$. If it happens, we may have to worry about domain-wall problem although the details may also depend on angular motion of $\Phi$. Thermal effect would be from the interaction of $\phi$ to RHNs which had been produced already from the partial decay of $\phi$. Those RHNs have energy of order of $m_{\phi}\approx\sqrt{3\lambda_{\phi}}\phi_{\times}$ at the epoch when $\phi\sim\phi_{\times}$ before either thermal or kinetic equilibrium is achieved. The number density is $n_{N_{i}}\approx 2{\rm B}_{i}\left(\Gamma^{\phi}/H_{\times}\right)\rho_{\phi}/m_{\phi}$, and their scattering cross-section to $\phi$ and axi-majorons is $\sigma=c_{\sigma}\times(\lambda_{\phi}^{4}/m_{\phi}^{2})$ with $c_{\sigma}=\mathcal{O}(0.1)$ being a numerical constant. Then, the ratio of the scattering rate $\Gamma_{s}$ to the expansion at the epoch, i.e., $H\sim H_{\times}\equiv H(\phi_{\times})$ with $H$ given in Eq. (3.62), is expected to be
$\frac{\Gamma_{s}}{H_{\times}}\approx\frac{3c_{\sigma}{\rm B}_{i}\lambda_{\phi}^{4}}{32\pi}\left(\frac{M_{\rm P}}{\phi_{0}}\right)^{2}.$ (3.68)
Hence, thermal effect can be ignored as long as the non-minimal gravitational coupling $\xi_{\phi}$ is constrained as
$\xi_{\phi}<\frac{350}{{\rm B}_{i}^{1/8}}\times\left(\frac{\phi_{0}}{10^{9}{\rm GeV}}\right)^{1/4}\ .$ (3.69)
For simplicity, we restrict ourselves to this case in the following discussion.
It should be noted that, even though thermal effect is negligible, domain-wall problem can be avoided only if $\Phi$ has angular motion which is large enough to overcome its quantum fluctuations. Otherwise, $\Phi$ would spread over the whole $2\pi$ angle, causing the domain-wall problem even though it may not get through a phase transition from the origin.
Now, we pay attention to the case of $\phi_{*}<\phi_{\times}$ first, which corresponds to
$\xi_{\phi}<100\times\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{1/2}\ ,$ (3.70)
and will comment on the opposite case (i.e., $\phi_{*}\gtrsim\phi_{\times}$) later. Once $\phi$ decays at the epoch of $H_{*}$, the universe will be filled with axi-majorons and RHNs as a comparable or sub-dominant component. Depending on interaction rates, they may or may not reach thermal or kinetic equilibrium among themselves before the standard thermal bath is established from decays of RHNs. Initially, the production of $\phi$ via inverse decay is kinematically forbidden. Hence, thermalization before the decay of RHNs would be driven by interactions of axi-majoron kinetic term, $y_{N_{i}}$, and $y_{\nu_{i}}$. The $\phi$-mediating self-interaction cross-section of axi-majorons is given by
$\sigma_{a,s}=\frac{3}{32\pi s}\left(\frac{m_{\phi}}{\phi_{0}}\right)^{4}\ ,$ (3.71)
where $s=E_{\rm CM}^{2}$ with $E_{\rm CM}$ the total energy at the center of momentum(CM) frame. Then the scattering rate is given by
$\Gamma_{a,s}=\frac{\lambda_{\phi}^{3}E_{a}}{128\pi^{3}}\left(\frac{M_{\rm P}}{\phi_{0}}\right)^{2}\ ,$ (3.72)
where $E_{a}$ is the energy of axi-majoron in the CM frame. The expansion rate of the universe during radiation-domination era scales as
$H(E_{a})=H_{*}\left(\frac{2E_{a}}{m_{\phi}}\right)^{2}\ .$ (3.73)
From the condition, $H\lesssim\Gamma_{a,s}/2$ for equilibration, the expansion rate at the epoch of axi-majoron kinetic-equilibrium among themselves is then given by
$H_{a,\rm equil}$ $=$ $H_{*}\times\left[\frac{\sqrt{3}\lambda_{\phi}}{8\pi}\left(\frac{M_{\rm P}}{\phi_{0}}\right)\right]^{4}$ (3.74)
$\simeq$ $H_{*}\times\left(\frac{\xi_{\phi}}{10^{2}}\right)^{8}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)^{4}\ .$
In the case of RHNs, the Yukawa-like interaction $y_{N}$ in Eq. (2.5) induces axi-majoron interactions to a Majorana field $N$,
$\mathcal{L}\supset-i\frac{y_{N}}{2\sqrt{2}}a_{\phi}\overline{N}\gamma_{5}N+\frac{y_{N}}{4\sqrt{2}\phi_{0}}a_{\phi}^{2}\overline{N}N+{\rm h.c.}\ ,$ (3.75)
where $N=\left(\nu_{R},\nu_{R}^{c}\right)^{T}$ is a four-component Majorana field with the flavor index suppressed, and only relevant leading order terms are shown. When the condition in Eq. (3.56) is satisfied for the RHN-channels of the decay of the inflaton, the cross-sections of the various interaction between the axi-majorons and the RHNs are smaller than the one for axi-majoron self-interactions by a factor of $\mathcal{O}(0.1-1)\times b_{i}^{2}$. On the other hand, from the ratio of $y_{\nu_{i}}$ to $y_{N_{i}}$, which is given by
$\frac{y_{\nu_{i}}^{4}}{y_{N_{i}}^{4}}=\left[\frac{m_{\nu}\phi_{0}}{v_{h}^{2}\sqrt{b_{i}\lambda_{\phi}}}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\right]^{2}\approx\frac{10^{3}}{b_{i}\xi_{\phi}^{2}}\left[\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\frac{m_{\nu_{i}}}{m_{\nu}}\right]^{2}\ ,$ (3.76)
one may think that the interaction $y_{\nu_{i}}$ can cause much more efficient scatterings and annihilations of $N_{i}$s to SM particles. However, the decay rate of a RHN mass eigenstate $N_{i}$ is given by (see Appendix C)
$\Gamma_{N_{i}}\simeq\frac{y_{\nu_{i}}^{2}m_{N_{i}}}{8\pi}=\frac{m_{\nu_{i}}}{4\pi}\left(\frac{m_{N_{i}}}{v_{h}}\right)^{2}=\frac{b_{i}\lambda_{\phi}m_{\nu_{i}}}{4\pi}\left(\frac{\phi_{0}}{v_{h}}\right)^{2}\ ,$ (3.77)
where $m_{N_{i}}\gg m_{h}$ was assumed. Also, if the left-handed neutrino mass-eigenstates have the normal hierarchy [13] for their masses as an example, $m_{\nu_{i}}\approx m_{\nu}$ is required for heavy RHN mass-eigenstates $N_{2,3}$. Then, for $\xi_{\phi}$ in the range we are interested in, with $\phi_{0}\sim\phi_{0}^{\rm ref}$, heavy RHN mass-eigenstates decay well before their self-interactions mediated by either axi-majorons or SM particles become relevant. The lightest RHNs can also be decoupled from axi-majorons and SM particles before their decay if $\xi_{\phi}\ll\mathcal{O}(10^{2})$ even though $m_{\nu_{1}}\ll m_{\nu}$ (see Eqs. (3.74) and (3.76)).
Once produced, depending on parameters, axi-majorons and RHNs may or may not evolve independently until the decay of the latter. If a RHN mass-eigenstate evolves independently at least until its decay, it becomes non-relativistic when the expansion rate becomes
$H_{i,\rm non}=H_{*}\left(\frac{2m_{N_{i}}}{m_{\phi}}\right)^{2}=2b_{i}H_{*}\ .$ (3.78)
It gives
$\frac{\Gamma_{N_{i}}}{H_{i,\rm non}}=\frac{3}{\sqrt{2\lambda_{\phi}}}\frac{m_{\nu_{i}}\phi_{0}}{v_{h}^{2}}\simeq\frac{66}{\xi_{\phi}}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.79)
Also, if stable enough, it would start dominating the universe when the momentum of axi-majoron becomes $p_{a}=p_{a,\rm eq}^{i}$ with
$p_{a,\rm eq}^{i}$ $=$ $2{\rm B}_{i}m_{i}\ .$ (3.80)
The expansion rate at this epoch is
$H_{i,\rm eq}=\left(\frac{p_{a,\rm eq}^{i}}{p_{a,*}}\right)^{2}H_{*}=8{\rm B}_{i}^{2}b_{i}H_{*}\ ,$ (3.81)
and one finds
$\frac{\Gamma_{N_{i}}}{H_{i,\rm eq}}=\frac{3}{4\sqrt{2\lambda_{\phi}}{\rm B}_{i}^{2}}\left(\frac{m_{\nu_{i}}\phi_{0}}{v_{h}^{2}}\right)\approx\frac{17}{{\rm B}_{i}^{2}\xi_{\phi}}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.82)
If one of RHN mass-eigenstates is responsible for eventual recovering the SM thermal bath, it would have to satisfy $\Gamma_{N_{i}}/H_{i,\rm eq}<0.1\times\left(\Delta N_{\rm eff}^{\rm obs}/0.5\right)^{3/2}$ with $\Delta N_{\rm eff}^{\rm obs}$ being the observational bound on the extra radiation (see Sectoin 5). Combined with the condition in Eq. (3.69), the branching ratio ${\rm B}_{i}$ is then constrained as
${\rm B}_{i}\gtrsim 2.0\times 10^{-2}\left(\frac{0.5}{\Delta N_{\rm eff}^{\rm obs}}\right)^{4/5}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)^{8/15}\left(\frac{\phi_{0}}{10^{9}{\rm GeV}}\right)^{2/5}\ .$ (3.83)
Hence, except the case of $\phi\sim 10^{9}{\rm GeV}$ with ${\rm B}_{i}=\mathcal{O}(0.1)$, typically heavy states are expected to decay around or before the epoch they become non-relativistic, and never dominate the universe. The standard thermal bath in this case can be recovered by the decay of the lightest RHN mass-eigenstate $N_{1}$ if $m_{\nu_{1}}$(or equivalently $y_{\nu_{1}}$) is sufficiently smaller than $m_{\nu}$.
Note that, as inferred from Eq. (3.83), if $\Delta N_{\rm eff}^{\rm obs}=\mathcal{O}(0.1)$ with $\phi_{0}\ll\phi_{0}^{\rm ref}$, heavy RHN states can decay while they dominate the universe, and recover the standard thermal bath by themselves. The lightest RHN $N_{1}$ in such a case can become the main component of dark matter, having self-interactions mediated by a very light axi-majoron. Although $N_{1}$ should have essentially zero Yukawa coupling $y_{\nu_{1}}$, in principle it is a possibility one can consider. We do not discuss this case further in this work, though.
If heavy RHN states decay before or around the epoch they become non-relativistic, their decay rate is suppressed by the boosting effect. For simplicity, we may take ${\rm B}_{N}\equiv{\rm B}_{3}\approx{\rm B}_{2}$, $\Gamma_{N}\equiv\Gamma_{N_{3}}\approx\Gamma_{N_{2}}$, $m_{\nu_{N}}\equiv m_{\nu_{3}}\approx m_{\nu_{2}}$, $m_{N}\equiv m_{N_{3}}\approx m_{N_{2}}$. Then, the expansion rate at the epoch of the decays of heavy RHNs is estimated as
$H_{N}=\frac{f_{N}(\xi_{\phi},\phi_{0})}{2}\Gamma_{N}\ ,$ (3.84)
where
$f_{N}(\xi_{\phi},\phi_{0})\equiv\left[\frac{H_{*}}{\Gamma_{N}/2}\left(\frac{m_{N}}{m_{\phi}}\right)^{2}\right]^{1/3}=\left[\frac{2\sqrt{2}\lambda_{\phi}^{1/2}}{3}\left(\frac{v_{h}^{2}}{m_{\nu_{N}}\phi_{0}}\right)\right]^{1/3}\simeq 0.3\xi_{\phi}^{1/3}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)^{1/3}\ ,$ (3.85)
with Eqs. (3.67) and (3.77) used for the last approximation. Note that Eq. (3.85) is valid only for $f_{N}\leq 1$. The temperature of the SM thermal bath established due to their decay is then given by
$T_{\gamma,N}$$\approx$ $\left(\frac{3\times 45{\rm B}_{N}}{\pi^{2}g_{*}(T_{\gamma,N})}\right)^{1/4}\sqrt{f_{N}\Gamma_{N}M_{\rm P}}$ (3.86)
$\simeq$ $2.4\times f_{N}^{1/2}\left(\frac{{\rm B}_{N}}{10^{-2}}\right)^{1/4}\left(\frac{m_{\nu_{N}}}{m_{\nu}}\right)^{1/2}m_{N}\ ,$
where we used Eq. (3.57) with $g_{*}(T_{\gamma,N})=106.75$ in the second line. Hence, unless ${\rm B}_{N}\ll 10^{-4}$, the heavy states $N_{2,3}$ can be thermalized with the SM thermal bath via inverse decay processes. Note that $m_{\phi}/T_{\gamma,N}\simeq 6.7\times f_{N}^{-1/2}\left(10^{-2}/b_{N}\right)^{3/4}$ with $b_{N}\equiv b_{3}\approx b_{2}$. Hence, if $b_{N}=\mathcal{O}(0.01-0.1)$, the inverse decay, $N+N\to\phi$, can also be efficient enough to thermalize $\phi$-particles. But the impact of the process is minor in our discussion. Axi-majorons are not thermalized even in this case, since the interaction is suppressed at least by $\mathcal{O}(T_{\gamma,N}^{2}/\phi_{0}^{2})$. There would be thermal contribution to the energy of axi-majorons from partial decays of thermalized $\phi$-particles, but it is subdominant.
4 Leptogenesis
Thanks to the symmetry-breaking $\xi_{a}$-term in Eq. (2.7), the inflation along $\Phi$-direction can generate angular motion at the end of inflation, as long as $\theta_{\rm ini}$ is not very close to zero and the mass scale along the angular direction is somewhat smaller but not extremely smaller than the expansion rate during inflation. Since $U(1)_{\rm PQ}$ is conserved in ${\tilde{V}}$, the angular momentum of $\Phi$ at the end of inflation can be conserved for $\phi\ll M_{\rm P}$, and corresponds to the Peccei-Quinn number asymmetry stored in $\Phi$:
$\Delta n_{\Phi}=i\left(\dot{\Phi}^{\dagger}\Phi-\Phi^{\dagger}\dot{\Phi}\right)=\dot{\theta}\phi^{2}=\dot{\phi}_{r}\phi_{i}-\phi_{r}\dot{\phi}_{i}\ ,$ (4.1)
with the PQ-charge of $\Phi$ normalized to be unity.
Here is a qualitative idea of how large PQ-number asymmetry can be generated in our model. At the end of inflation, when $m_{a,e}^{2}/3H_{e}^{2}\lesssim\mathcal{O}(10^{-2})$ to avoid too fast angular motion toward the angular minimum position during inflation, it is expected that
$\dot{\theta}_{e}\sim\theta_{\rm ini}\left(\frac{m_{a,e}^{2}}{H_{e}^{2}}\right)H_{e}\simeq\frac{\sqrt{3}}{3}\theta_{\rm ini}\alpha\xi_{\phi}\lambda_{\phi}^{\rm ref}M_{\rm P}\ ,$ (4.2)
where Eq. (2.41) and
$H_{e}\approx\frac{\lambda_{\phi}^{1/2}\phi_{e}^{2}}{2\sqrt{3}M_{\rm P}}\simeq\frac{\sqrt{\lambda_{\phi}^{\rm ref}}}{2\sqrt{3}}M_{\rm P}\ ,$ (4.3)
with $\phi_{e}\simeq M_{\rm P}/\sqrt{\xi_{\phi}}$ [38] was used in the last approximation. Also, for convenience, one may define the effective temperature of the radiation-like energy component at the end of inflation as
$T_{e}\approx\left(\frac{\pi^{2}}{90}g_{*}^{\Phi}\right)^{-1/4}\sqrt{H_{e}M_{\rm P}}=\left(\frac{2\pi^{2}g_{*}^{\Phi}}{15}\right)^{-1/4}\lambda_{\phi}^{1/4}\phi_{e}\ ,$ (4.4)
where $g_{*}^{\Phi}=2$ is the degrees of freedom of $\Phi$. Then, in terms of the yield defined as the ratio of the PQ-number density to the entropy density, the PQ-asymmetry at the end of inflation can be express as
$Y_{\Phi,e}$ $=$ $\frac{3}{4}\frac{T_{e}\Delta n_{\rm PQ,e}}{\rho_{\phi,\rm e}}=\frac{3}{2\lambda_{\phi}}\frac{T_{e}\dot{\theta}_{e}}{\phi_{e}^{2}}$ (4.5)
$\approx$ $2.5\left(\frac{\alpha}{10^{-2}}\right)\times\theta_{\rm ini}\ .$
It seems that $Y_{\Phi,e}$ depends only on $\alpha$, when $\lambda_{\phi}$ and $\xi_{\phi}$ satisfy the relation in Eq. (3.51) to provide inflationary observables matching observations. However, it should be noted that Eq. (4.5) provides just a qualitative idea of the maximal possibility of $Y_{\Phi,e}$. The exact parametric dependence of $Y_{\Phi,e}$ is non-trivial, since it involves the motion of $\Phi$ at the end of inflation that is affected by the initial position of $\phi$, i.e., $\phi_{\rm ini}$. For later convenience, we may express $\dot{\theta}_{e}$ in terms of $Y_{\Phi,e}$ as
$\dot{\theta}_{e}=\frac{2}{3}\left(\frac{2\pi^{2}}{15}\right)^{1/4}\lambda_{\phi}^{3/4}\phi_{e}Y_{\Phi,e}\ .$ (4.6)
Fig. 8 shows numerical estimations of $Y_{\Phi}$ well after inflation as functions of $\theta_{\rm ini}$ for various values of $\alpha$.
Figure 8: The yield of the PQ-number asymmetry $Y_{\rm PQ}$ at $t_{f}=10^{4}H_{\rm ref}^{-1}$ as functions of $\theta_{\rm ini}$ for various $\alpha$ with $\phi_{\rm ini}/M_{\rm P}=10$(left), $20$(right). Inflation ends at $t_{e}=\mathcal{O}(100)H_{\rm ref}^{-1}$ depending on initial position of the inflaton and $\alpha$.
The very different behaviors of the $\alpha=10^{-2}$ case in the left and right panels of the figure are due the fact that the angular motion during inflation is sizable for a large $\alpha$ so that, if $\phi_{\rm ini}$ is quite large, the field configuration can move the direction of the minimum along angular direction first, and towards the origin, afterward. From the figure, one can see that, for $\alpha\lesssim 10^{-2}$ it is possible to obtain a quite large amount of asymmetry at least for $\theta_{\rm ini}=\mathcal{O}(0.1-1)$ which may be quite natural.
The eventual decays of $\Phi$ field transfer the PQ-asymmetry to the asymmetry of RHN-helicity states via the interaction $y_{N}$ responsible for generating the majorana mass of RHNs. One may assign 2-units of lepton number to the Peccei-Quinn field, and one-unit to the right-handed neutrino $\nu_{R}$. Then, both interactions $y_{N}$ and $y_{\nu}$ respects not only $U(1)_{\rm PQ}$ but also $U(1)_{B-L}$, and the PQ-number asymmetry of $\Phi$ field can be regarded as the lepton number asymmetry as $\Delta n_{L}^{\Phi}=2\Delta n_{\Phi}$. In the broken phase of $U(1)_{\rm PQ}$-symmetry, the non-zero majorana mass-term breaks the lepton number conservation. However, as eigenstates of the equation of motion, helicity states are not affected by the majorana mass-term and the asymmetry between the helicity states of RHNs is not changed. The Yukawa interaction $y_{\nu}$ violates the helicity conservation, but the effect depends on the speed of RHNs. The branching ratio of the channel, $N_{i}^{+}\to\ell+H$, is $\left(1+\beta_{i}\right)/2$ whereas that of the channel, $N_{i}^{+}\to\overline{\ell}+\overline{H}$, is $\left(1-\beta_{i}\right)/2$ with $\beta_{i}$ being the speed of $N_{i}$. Hence, the helicity asymmetry of $N_{i}$ induced from the decay of $\Phi$-field is only partially transferred to the visible sector by the suppression factor $\beta_{i}$. As a result, if $N_{i}$ is never equilibrated, its decay contributes to the lepton number asymmetry of the visible sector as [60]
$\Delta n_{L}^{i}=\Delta n_{\rm hel}^{i}\times\beta_{i}$ (4.7)
where $\Delta n_{\rm hel}^{i}$ denotes the asymmetry between $N_{i}^{+}$ and $N_{i}^{-}$.
The late time asymmetry in the visible sector depends not only on the speeds of RHNs but also on the background temperature of the SM thermal bath when RHNs decay. Depending on parameters causing different situations, there are several possibilities for generating baryon number asymmetry of the universe: Affleck-Dine (AD) baryogenesis [27] through the decay chain such as $\Phi\to\nu_{R}\to\ell$ with RHNs out of equilibrium, spontaneous leptogenesis thanks to the motion of the axi-majoron field with thermal RHNs [61], and resonant leptogenesis [28] from decays of non-thermal RHNs.
4.1 Peccei-Quinn Affleck-Dine leptogenesis
We consider the case in which the decay of $N_{i}$ happen in non-relativistic regime and the background temperature is much lower than $m_{N_{i}}$ so that washing-out processes are not active. Looking at Eqs. (3.82), (3.83) and (3.86), we see that this is relevant only for the lightest RHN $N_{1}$ for which $m_{\nu_{1}}$ can be much smaller than $m_{\nu}$, and $N_{1}$ should be responsible for recovering the late time SM thermal bath.
A part of the PQ-number asymmetry can be transferred to the visible sector through the helicity states of $N_{1}$, while the contributions from heavier states are washed to a much smaller value. It is possible for the heavy states to decay when they just become non-relativistic while producing SM particles with a temperature $T\sim m_{N}\gtrsim m_{N_{1}}$. For a diagonal mass matrix of RHNs, which we are assuming in this work, the scattering and inverse process to thermalize $N_{1}$ is efficient only when $T\sim m_{N_{1}}$ with $T$ being the temperature in the standard radiation-dominated universe. In such a thermal bath, the scattering rate of $N_{1}$ to SM particles when $T\gtrsim m_{N_{1}}$ is expected to be $\Gamma_{1,s}=c_{1,s}y_{\nu_{1}}^{2}T$ with $c_{1,s}=\mathcal{O}(10^{-2})$ [60] being a numerical factor including the effect of SM couplings and the number of allowed channels. Note that the expansion rate at $T=m_{N_{1}}$ is given by
$H(T=m_{N_{1}})=H_{\rm SM}(T=m_{N_{1}})/\sqrt{2{\rm B}_{N}}$
with $H_{\rm SM}(T)$ being the expansion rate in the standard cosmology in a radiation-dominated universe with temperature $T$. The ratio of $\Gamma_{1,s}$ to the expansion rate is then found as
$\frac{\Gamma_{1,s}}{H(T=m_{N_{1}})}$ $=$ $2\left(\frac{90}{\pi^{2}g_{*}(T=m_{N_{1}})}\right)^{1/2}c_{1,s}\sqrt{2{\rm B}_{N}}\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}$ (4.8)
$\approx$ $5.2\times\left(\frac{c_{1,s}}{10^{-2}}\right)\left(\frac{{\rm B}_{N}}{0.1}\right)^{1/2}\left(\frac{m_{\nu_{1}}}{m_{\nu}}\right)\ ,$
where $g_{*}(T=m_{N_{1}})=106.75$ was used in the second line. Hence, for $m_{\nu_{1}}$ smaller than $m_{\nu}$ by a couple of orders of magnitude that may be necessary in any case for $N_{1}$ to recover the SM thermal bath finally (see Eq. (3.82)), the scattering of $N_{1}$ can be inefficient. Then, even though there would be thermal contribution to the abundance of $N_{1}^{\pm}$s as a subdominant portion, once produced in the decay of the inflaton, $N_{1}$ can be never thermalized with the SM particles.
The details of the transfer processes of the PQ-number asymmetry to lepton-number asymmetry depend on whether $\phi_{*}$ is larger or smaller than $\phi_{\times}$. If $\phi_{*}<\phi_{\times}$, contrary to the conventional scenarios of AD-mechanism in which a AD-field carrying the asymmetry decay always in a symmetric phase, $\Phi$ falls into the vicinity of the true vacuum position before it finally decays. At the broken phase, $\Phi$ is decomposed into radial and angular (axi-majoron) modes. The radial mode does not carry any asymmetry, but it will heat up the universe by its decay. Hence, AD-mechanism is expected to be relevant only when the oscillation amplitude of $\Phi$ is larger than $\phi_{\times}$ at which the energy of field is the same as the one at the origin. The precise criterion to distinguish those two regime is not clear. For simplicity, we assume that it happens at the epoch of $H_{\times}$ and include the uncertainty as a suppression parameter $c_{\rm AD}$. Then, the helicity asymmetry associated with $N_{1}$ from the partial decay of $\Phi$ at the epoch is given by
$\left(\Delta n_{\rm hel}^{1}\right)_{\times}\approx\frac{4c_{\rm AD}{\rm B}_{1}}{5}\frac{\Gamma^{\phi}}{H_{\times}}\left(\Delta n_{\Phi}\right)_{\times}\ .$ (4.9)
It has a suppression factor
$\Delta_{\rm AD}\equiv\left(c_{\rm AD}\frac{2\Gamma^{\phi}}{5H_{\times}}\right)\ ,$ (4.10)
relative to the expectation in conventional scenarios. For the decoupled $N_{1}$, the speed $\beta_{1}$ at its decay is given by
$\beta_{1}=\left(\frac{H_{1,\rm eq}}{H_{1,\rm non}}\right)^{1/2}\left(\frac{H_{\rm d}}{H_{1,\rm eq}}\right)^{2/3}=\left({\rm B}_{1}\lambda_{\phi}\right)^{-1/3}\left(\frac{m_{\nu_{1}}\phi_{0}}{v_{h}^{2}}\right)^{2/3}\ .$ (4.11)
Then, the late time asymmetry transferred through the decays of $N_{1}$s is found to be
$Y_{B,{\rm AD}}^{<}$ $=$ $-\Delta_{\rm AD}\times\beta_{1}\times\frac{12}{37}\left(\frac{\sqrt{3}}{8\pi}\right)^{1/2}\left(\frac{g_{*}^{\Phi}}{g_{*}(T_{\rm d})}\right)^{1/4}\lambda_{\phi}^{1/4}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}\right)^{1/2}Y_{\Phi,e}$
$\simeq$ $-1.7\times 10^{-10}\times\left(\frac{c_{\rm AD}}{0.1}\right)\left(\frac{0.1}{{\rm B}_{1}}\right)^{1/3}\left(\frac{\xi_{\phi}}{10}\right)^{11/6}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{7/6}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)^{1/3}\left(\frac{Y_{\Phi,e}}{10^{-1}}\right)\ ,$
where we used $\Gamma^{\phi}\approx\Gamma_{a}^{\phi}$ in the second line. Since axi-majorons produced by misalignment are the only contribution to dark matter and kinetic misalignment does not work in our scenario (see Section 5), it is required to have $\phi_{0}\sim\phi_{0}^{\rm ref}$. Also, the validity of Eq. (4.1) is limited to the case of $\phi_{*}\lesssim\phi_{\times}$ which puts a limit $\xi_{\phi}\lesssim 100$ (see Eq. (3.70)). Moreover, the constraint of $\Gamma_{N_{1}}\ll H_{1,\rm eq}$ (see Eq. (3.82)) requires ${\rm B}_{1}\gtrsim\left(m_{\nu_{1}}/m_{\nu}\right)^{1/2}$. Therefore, if $c_{\rm AD}$ is not much smaller than unity, $Y_{\Phi,e}$ should be smaller than unity by a couple of orders of magnitude in order to match observation.
If $\phi_{*}>\phi_{\times}$, $\Phi$ field decays before falling into the region of the true vacuum position. As shown in Eq. (3.63), this case would be possible if $\xi_{\phi}\gtrsim\mathcal{O}(10^{2})$ for $\phi_{0}\sim\phi_{0}^{\rm ref}$. Thermalization of $N_{1}$ does not occur as long as $\xi_{\phi}\sim\mathcal{O}(10^{2})$ while keeping $\phi_{0}\approx\phi_{0}^{\rm ref}$. The late time baryon number asymmetry in this case is given by
$Y_{B,\rm AD}^{>}$ $=$ $-\beta_{1}\times\frac{12}{37}\left(\frac{3\sqrt{3}}{2\pi}\right)^{1/2}\left(\frac{g_{*}^{\Phi}}{g_{*}(T_{\rm d})}\right)^{1/4}\lambda_{\phi}^{1/4}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}\right)^{1/2}Y_{\Phi,e}$ (4.13)
$\simeq$ $-7.0\times 10^{-11}\times\left(\frac{0.1}{{\rm B}_{1}}\right)^{1/3}\left(\frac{200}{\xi_{\phi}}\right)^{1/6}\left(\frac{10^{3}m_{\nu_{1}}}{m_{\nu}}\right)^{7/6}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{2/3}\left(\frac{Y_{\Phi,e}}{10^{-6}}\right)$
Compared to Eq. (4.1), Eq. (4.13) does not have the suppression factor $\Delta_{\rm AD}$, and a smaller $Y_{\Phi,e}$ is required for a given set of parameters.
Figure 9: Parameter for a baryon number asymmetry (red solid line), matching observations $Y_{B}^{\rm obs}\simeq 9\times 10^{-11}$ [62], from the contribution of the lightest RHN via AD-mechanism as a function of $m_{\nu_{1}}$ and $b_{1}$ for $b_{N}=0.1$, and $\phi_{0}=\phi_{0}^{\rm ref}$. For a smaller $Y_{\Phi,e}$, the red line is pushed up. The blue shaded region may be excluded due to too much dark radiation. Blue liens corresponds to $\Gamma_{1}/H_{\rm eq}=10^{-3}$(dotted blue), $10^{-2}$(dashed blue), and $0.1$(blue). Green lines are for $T_{\gamma,N_{1}}/m_{N_{1}}=0.05$(dashed green) and $0.1$(solid blue). Left: $Y_{B,\rm AD}^{<}$ with $Y_{\Phi,e}=10^{-1}$ in Eq. (4.1) for $\xi_{\phi}=10$ leading to $\phi_{*}/\phi_{\times}=9.1\times 10^{-3}$. Right: $Y_{B,\rm AD}^{>}$ with $Y_{\Phi,e}=10^{-6}$ in Eq. (4.13) for $\xi_{\phi}=200$ leading to $\phi_{*}/\phi_{\times}=3.6$.
Fig. 9 shows the parameter space for $Y_{B,\rm AD}$ matching observations. As shown in the figure, we can obtain the right amount of baryon asymmetry if $m_{\nu_{1}}$ is smaller than $m_{\nu}$ by several orders of magnitude while a sizable extra-radiation can also be obtained from hot axi-majorons (see Section 5).
A remark is in order before closing this subsection. The initial asymmetry stored in $\Phi$ is only partially transferred to the visible sector with the help of RHNs. The remaining part of the asymmetry is kept as the background motion of axi-majoron. Since $U(1)_{\rm PQ}$ symmetry is broken only spontaneously in the matter sector, the amount of the net PQ-number asymmetry in a co-moving volume after inflation should be conserved. This means that the loss of asymmetry in the visible sector should appear as the motion of axi-majoron, but has minor impact on the overall homogeneous motion. Meanwhile, if the background axi-majoron motion is large enough, it can become the main source of baryogenesis via spontaneous leptogenesis as discussed in Ref. [61] for example. We discuss this possibility in the next subsection.
4.2 Spontaneous leptogenesis via axi-majoron
If RHNs are in thermal bath while they are coupled to the derivative of the axi-majoron field, the idea of spontaneous baryogenesis [40] can be applied. In Ref. [61], it was shown that in the presence of a non-zero $\dot{\theta}$, the derivative couplings of the majoron(the axi-majoron in our scenario) to fermion currents can act as a source term for generating non-zero lepton number asymmetry via inverse decay processes ($\ell+H\to N_{i}^{\pm}$) in equilibrium through the Yukawa interaction $y_{\nu}$. In our scenario, due to its nature of sharing the properties of axion and majoron, axi-majoron contributes to QCD sphaleron [63] as well as electroweak sphaleron processes, and Boltzmann equations are modified similarly to the case discussed in Ref. [39]. However, once $L$-number violation is frozen when the inverse decay to $N_{i}^{\pm}$ is closed, $B-L$ is also frozen. Also, $B+L$ is eventually driven to cancel axi-majoron contribution in the electroweak sphaleron contribution in Boltzmann equations, and $\dot{\theta}/T$ at the freeze-out epoch of the sphalerons is negligible in our scenario as will be clear in the subsequent discussion. Hence, the eventual baryon and lepton number asymmetries are determined by the non-zero $B-L$ number which is determined in a way similar to Ref. [61]. The effect of QCD/electroweak sphalerons on the results is minor, and we will ignore such a difference since it is easily compensated by other parameters.
If $N_{1}$ decays out of equilibrium and is never thermalized that is the case discussed in the previous section, only heavy RHNs are relevant for spontaneous leptogenesis. The temperature of the SM thermal bath established by their decays is given in Eq. (3.86), and one finds
$H(T_{\gamma}=m_{N_{i}})=\frac{\Gamma_{N_{i}}}{2}\left(\frac{m_{N_{i}}}{T_{\gamma,N}}\right)^{2}\ ,$ (4.14)
leading to
$\eta_{i}$$\equiv$ $\frac{\Gamma_{N_{i}}}{H(T_{\gamma}=m_{N_{i}})}\simeq 66\times{\rm B}_{N}^{1/2}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\ .$ (4.15)
Therefore, the wash-in effect is strong enough as long as ${\rm B}_{N}\gtrsim\mathcal{O}(10^{-4})$. The processes will be shut down as the background temperature becomes sufficiently smaller than the mass of relevant RHN mass-eigenstates. The expansion rate at that epoch is found to be [61]
$H_{\rm fo}=c_{\rm inv}\Gamma_{N}\ ,$ (4.16)
with $c_{\rm inv}\simeq 2.7\times 10^{-3}$. In the standard thermal bath, the freeze-out temperature is then given by
$T_{\gamma,{\rm fo}}=\frac{m_{N}}{x_{\rm fo}}=\frac{\sqrt{b_{N}\lambda_{\phi}}}{x_{\rm fo}}\phi_{0}\simeq 2.6\times 10^{6}{\rm GeV}\times b_{N}^{1/2}\xi_{\phi}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ ,$ (4.17)
with $x_{\rm fo}\approx 10$. In our scenario $H_{\rm fo}$ is larger than the one related to $T_{\gamma,\rm fo}$ in the standard cosmology, leading to a smaller $x_{\rm fo}$. However, such a dependence is mainly logarithmic on ${\rm B}_{N}$, and minor as long ${\rm B}_{N}=\mathcal{O}(0.1)$ which we assume in this work. In this reason, we ignore such a dependence.
The baryon and lepton number asymmetry at the freeze-out in the singlet majoron scenario is found as [61]
$\left(n_{B,L}\right)_{\rm fo}=\frac{c_{B,L}}{6}\left(\dot{\theta}T_{\gamma}^{2}\right)_{\rm fo}\ ,$ (4.18)
where $c_{B}$ and $c_{L}$ are model-dependent numerical coefficients of order unity. Angular velocity $\dot{\theta}$ after inflation scales as
$\dot{\theta}\propto\left\{\begin{array}[]{ll}a^{-1}&:\ \phi\gtrsim\phi_{\times}\ ,\\ a^{-3}&:\ \phi\lesssim\phi_{\times}\ .\end{array}\right.$ (4.19)
Hence, $\dot{\theta}T_{\gamma}^{2}\propto a^{-5}$ for ${\tilde{\phi}}_{*}\ll\phi_{0}$. Since the standard thermal bath is established by the decay of the lightest RHN $N_{1}$ which should dominate the universe for a while, there would be a dilution of pre-existing baryon and lepton number asymmetries. If $N_{1}$ decays well before the epoch of the EWPT, the late time baryon number asymmetry through the spontaneous baryogenesis mechanism is given by [64]
$Y_{B,\rm sb}=\frac{12}{37}Y_{B-L}^{\rm d}$ (4.20)
where $Y_{B-L}^{\rm d}$ is the yield of $B-L$ charge at the epoch of $N_{1}$’s decay. One finds
$Y_{B-L}^{\rm d}$ $=$ $Y_{B-L}^{\rm fo}\left(\frac{s_{\rm fo}}{s_{\rm d}}\right)\left(\frac{a_{\rm fo}}{a_{\rm d}}\right)^{3}$ (4.21)
$=$ $4{\rm B}_{N}x_{\rm fo}Y_{B-L}^{\rm fo}\left(\frac{T_{\rm d}}{m_{N}}\right)\left(\frac{H_{\rm fo}}{H_{\rm eq}}\right)^{1/2}$
where $T_{\rm d}$ is the decay temperature of $N_{1}$, and $H_{\rm eq}\equiv H_{1,\rm eq}$ given by Eq. (3.81). Note that Eq. (4.21) is valid for $H_{\rm fo}/H_{\rm eq}\geq 1$. The $B-L$ asymmetry at freeze-out is given by
$Y_{B-L}^{\rm fo}=-\frac{45c_{B-L}}{12\pi^{2}g_{*S}(T_{\gamma,\rm fo})}\left(\frac{\dot{\theta}}{T}\right)_{\rm fo}\ ,$ (4.22)
where $c_{B-L}=\mathcal{O}(1)$ [61]. In the case of $\phi_{*}<\phi_{\times}$, the velocity of axi-majoron at the epoch of $H_{\rm fo}$ is given by
$\dot{\theta}_{\rm fo}\approx\dot{\theta}_{e}\left(\frac{\phi_{\times}}{\phi_{e}}\right)\left(\frac{H_{*}}{H_{\times}}\right)^{2}\left(\frac{H_{\rm fo}}{H_{*}}\right)^{3/2}\ ,$ (4.23)
with
$H_{\times}\equiv\frac{\lambda_{\phi}^{1/2}\phi_{\times}^{2}}{2\sqrt{3}M_{\rm P}}\ .$ (4.24)
Also, the decay temperature of $N_{1}$ is found as
$\frac{T_{\rm d}}{m_{N}}=\left(\frac{5}{2\pi^{4}g_{*}(T_{\rm d})}\right)^{1/4}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}\right)^{1/2}\left(\frac{b_{1}}{b_{N}}\right)^{1/2},$ (4.25)
and
$\left(\frac{H_{\rm fo}}{H_{\rm eq}}\right)^{1/2}$ $=$ $c_{\rm inv}^{1/2}\left(\frac{b_{N}m_{\nu_{N}}}{b_{1}m_{\nu_{1}}}\right)^{1/2}\left(\frac{\Gamma_{N_{1}}}{H_{\rm eq}}\right)^{1/2}$ (4.26)
$=$ $\left[\frac{3c_{\rm inv}}{4\sqrt{2}}\frac{b_{N}}{b_{1}{\rm B}_{1}^{2}\lambda_{\phi}^{1/2}}\left(\frac{m_{\nu_{N}}\phi_{0}}{v_{h}^{2}}\right)\right]^{1/2}\ .$
Combining all, we find the late-time baryon number asymmetry for $\phi_{*}<\phi_{\times}$ as
$Y_{B,\rm sb}^{<}$$\approx$ $-\frac{2c_{B-L}c_{\rm inv}^{2}x_{\rm fo}^{2}}{\pi^{9/2}g_{*}(T_{\rm d})^{1/4}g_{*}(T_{\gamma,\rm fo})}\frac{b_{N}^{2}\lambda_{\phi}^{5/4}}{b_{1}}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}\right)^{1/2}\left(\frac{m_{\nu_{N}}M_{\rm P}}{v_{h}^{2}}\right)^{2}Y_{\Phi,e}$ (4.27)
$\simeq$ $-3.1\times 10^{-10}\left(\frac{0.05}{b_{1}}\right)\left(\frac{b_{N}}{0.15}\right)^{2}\left(\frac{\xi_{\phi}}{15}\right)^{5/2}\left(\frac{10^{3}m_{\nu_{1}}}{m_{\nu}}\right)^{1/2}Y_{\Phi,e}\ .$
where we used $c_{B-L}=2,\ g_{*}(T_{\rm d})=g_{*}(T_{\gamma,\rm fo})=106.75$, and $m_{\nu_{N}}=m_{\nu}$ in the second line. Note that, interestingly, $Y_{B}$ does not depend on the symmetry-breaking scale.
Figure 10: Parameter space of $H_{\rm fo}/H_{1,\rm eq}=1$(red line), $H_{\rm d}/H_{1,\rm eq}\leq 0.1$(blue solid line), $0.01$(blue dashed line) with $H_{\rm d}=(2/3)\Gamma_{N_{1}}$, and $\phi_{*}/\phi_{\times}=1$(green line) for $b_{1}=0.05,\ b_{2}=b_{3}=0.15$, and $m_{\nu_{N}}=m_{\nu}=10^{3}m_{\nu_{1}}$. The position of green line is nearly independent with branching ratios of $N_{i}$s. In the region left of the green line, since $\Gamma_{N_{i}}/H_{1,\rm eq}\propto b_{i}\phi_{0}/{\rm B}_{1}^{2}b_{1}\xi_{\phi}$, reducing $b_{2,3}$ shifts the blue line upward while red lines are also move upward due to the $b_{N}$-dependence of ${\rm B}_{1}$ even for a fixed $b_{1}$. A smaller $m_{\nu_{1}}$ also pushes up the blue lines.
In Fig. 10, we depicted parameter spaces for different cases. The condition $H_{\rm fo}/H_{\rm eq}\geq 1$ is satisfied in the region left of the red line. The blue region is nearly excluded due to too much dark radiation which is discussed in Section 5. However, the region can be pushed up by taking $m_{\nu_{1}}$ small enough. We also show in Fig. 11 the parameter space for $Y_{B,\rm sb}^{<}$ matching observations with $\Gamma_{1}/H_{\rm eq}\ll 1$ satisfied to recover the standard thermal bath. In the figure, as $\xi_{\phi}$ increases, the red line moves down or right while blue lines and shaded region move to opposite direction, as clear from Eqs. (3.82) and (4.27).
Figure 11: The late time baryon number asymmetry, $Y_{B,\rm sb}^{<}/Y_{\Phi,e}=Y_{B}^{\rm obs}$(red line), in Eq. (4.27) as a function of $m_{\nu_{1}}$ and $b_{1}$ for $b_{N}=0.15$, $\xi_{\phi}=15$, and $\phi_{0}=\phi_{0}^{\rm ref}$. The scheme of the blue region and the blue lines is the same as the one in Fig. 9. In the region left of the dashed green line, $H_{\rm fo}>H_{\rm eq}$ is satisfied.
As shown in Fig. 10, and inferred from Eq. (4.26), if $\xi_{\phi}\gtrsim\mathcal{O}(10^{2})$, the case of $H_{\rm fo}<H_{\rm eq}$ can also arise. In such case the heavy RHN states can also dominate the universe, and recover the standard thermal bath by their decay (see Eq. (3.82)). It allows the possibility of dark matter made of mainly the lightest RHNs when $\phi_{0}$ is not large enough to saturate the relic density by axi-majoron dark matter. Keeping this possibility in mind, we consider the case of $\xi_{\phi}\gtrsim\mathcal{O}(10^{2})$ with all RHNs having nearly same mass parameters and Yukawa couplings $y_{\nu}$. In regard of baryogenesis, this case is nearly the same as the case of the lightest stable and two nearly degenerate heavy RHN states, as long as the branching fraction to the lightest one is sufficiently small in order for it not to over-close the universe. Note that, as can be seen from Eq. (3.63), a large $\xi_{\phi}$ also leads to $\phi_{*}>\phi_{\times}$, i.e., the universe is dominated by a radiation-like energy of $\phi$-field at the epoch of $\phi$-particles’ decay. The expansion rate at the epoch of the inflaton’s decay in this case is then given by
$H_{*}=\frac{\lambda_{\phi}^{1/2}}{2\sqrt{3}}\frac{\phi_{*}^{2}}{M_{\rm P}}\ .$ (4.28)
Since RHNs having nearly the same decay rates are responsible for the final reheating,
$\frac{H_{\rm fo}}{H_{\rm d}}\approx\frac{3}{2}c_{\rm inv}\ ,$ (4.29)
where $H_{\rm d}\equiv(2/3)\Gamma_{N}$, and the late time baryon number asymmetry is given by
$Y_{B,\rm sb}^{>}=\frac{12}{37}Y_{B-L}^{\rm fo}$ (4.30)
for $N_{i}$s decaying well before the epoch of the EWPT. The velocity of axi-majoron at the epoch of $H_{*}$ is given by
$\dot{\theta}_{*}=\dot{\theta}_{e}\left(\frac{\phi_{0}}{\phi_{e}}\right)\left(\frac{\phi_{*}}{\phi_{0}}\right)^{3}\ ,$ (4.31)
and the freeze-out velocity of axi-majoron is given by
$\dot{\theta}_{\rm fo}\approx\dot{\theta}_{*}\left(\frac{H_{\rm eq}}{H_{*}}\right)^{3/2}\left(\frac{H_{\rm d}}{H_{\rm eq}}\right)^{2}\left(\frac{H_{\rm fo}}{H_{\rm d}}\right)^{3/2}\ ,$ (4.32)
where $H_{\rm eq}\equiv H_{3,\rm eq}$ given in Eq. (3.81). Combining all, in this case we find
$Y_{B,\rm sb}^{>}$$\approx$ $-\frac{c_{B-L}c_{\rm inv}^{3/2}x_{\rm fo}\lambda^{1/4}}{\pi^{5/2}g_{*}(T_{\gamma,\rm fo})}\left(\frac{b_{N}}{{\rm B}_{N}}\right)\left(\frac{m_{\nu_{N}}M_{\rm P}}{v_{h}^{2}}\right)^{2}\left(\frac{\phi_{0}}{M_{\rm P}}\right)Y_{\Phi,e}$ (4.33)
$\simeq$ $-1.6\times 10^{-10}\left(\frac{\xi_{\phi}}{10^{2}}\right)^{1/2}\left(\frac{\phi_{0}}{10^{9}{\rm GeV}}\right)Y_{\Phi,e}\ ,$
where we have used $b_{N}=0.1$, $m_{\nu_{1}}=m_{\nu_{N}}=m_{\nu}$ in the second line. The observed baryon asymmetry can be obtained by interplay between $\xi_{\phi}$, $\phi_{0}$, and $Y_{\Phi,e}$.
Figure 12: The late time baryon number asymmetry, $Y_{B}^{>}=Y_{B}^{\rm obs}$(red lines) in Eq. (4.33) as a function of $\xi_{\phi}$ and $\phi_{0}$ for $b_{1}=b_{2}=b_{3}=0.1$. Red lines correspond to the cases of $Y_{\Phi,e}=10^{-2}$(dotted), $10^{-1}$(dashed), and $1$(solid). Also, depicted are $\Gamma_{N,3}/H_{3,\rm eq}=10^{-3}$(dotted blue), $10^{-2}$(dashed blue), and $0.1$(blue). The blue shaded region may be excluded due to too much dark radiation. $H_{\rm fo}<H_{\rm d}$ is always satisfied.
Fig. 12 shows the parameter space of $Y_{B,\rm sb}^{>}$ matching observations as a function of $\xi_{\phi}$ and $\phi_{0}$. Note that $\phi_{0}\ll\phi_{0}^{\rm ref}$ is also allowed with a sizable amount of dark radiation from hot axi-majorons (see Section 5).
4.3 Resonant leptogenesis
If the inflaton does not carry enough amount of asymmetry, the baryon number asymmetry of the present universe should be obtained by decays of RHNs, following the usual way of out-of-equilibrium decay. Because of the tiny $y_{\nu}$ in our scenario, the so-called ’resonant leptogenesis’ [28] may be the only possibility in this case. The essential point of the mechanism is that the CP violation in the decay of $N_{i}$ mass-eigenstate can be given by
$\delta_{i}\approx{\tilde{\delta}}_{ij}\times\frac{\left(m_{i}^{2}-m_{j}^{2}\right)m_{i}\Gamma_{j}}{\left(m_{i}^{2}-m_{j}^{2}\right)^{2}+m_{i}^{2}\Gamma_{j}^{2}}\ \xrightarrow[]{\left(\Delta m\right)_{ij}\ll m_{i}}\ {\tilde{\delta}}_{ij}\times\frac{2\left(\Delta m\right)_{ij}m_{i}^{2}\Gamma_{j}}{4m_{i}^{2}\left(\Delta m\right)_{ij}^{2}+m_{i}^{2}\Gamma_{j}^{2}}\ ,$ (4.34)
where
${\tilde{\delta}}_{ij}\equiv\left[\frac{{\rm Im}\left(y_{\nu}^{\dagger}y_{\nu}\right)_{ij}^{2}}{\left(y_{\nu}^{\dagger}y_{\nu}\right)_{ii}\left(y_{\nu}^{\dagger}y_{\nu}\right)_{jj}}\right]$ (4.35)
and $\left(\Delta m\right)_{ij}\equiv|m_{i}-m_{j}|$. Hence, if $\left(\Delta m\right)_{ij}\sim\Gamma_{i}\simeq\Gamma_{j}$, it is possible to have $\delta_{i}\sim\mathcal{O}(1)$. We assume that two heavier RHN mass-eigenstates, $N_{2,3}$ are responsible for the resonant leptogenesis while the lightest recovers the standard thermal bath via its delayed decay (see Section 3.3). Lepton number asymmetry is produced well before the electroweak phase transition. It is diluted by the later entropy injection from the decay of the long-lived lightest RHN, $N_{1}$. The present baryon number asymmetry is then expected to be [28]
$Y_{B,\rm res}\sim\sum_{i=2,3}\frac{\delta_{i}}{1400\eta_{i}}\left(\frac{\Gamma_{1}}{H_{\rm eq}}\right)^{1/2}$ (4.36)
where the dilution factor $\eta_{i}$ is given in Eq. (4.15)
The mass splitting between $N_{2}$ and $N_{3}$ before their decay is mainly from 1-loop corrections  [^19] The SM thermal bath is initially established from the decay of those heavy RHNs. Thermal correction to the mass splitting of $N_{2}$ and $N_{3}$ before their decay is negligible.. It is given by [65]
$\left(\Delta m\right)=c_{m}\frac{y_{\nu}^{2}}{16\pi}m_{N}=\frac{c_{m}}{2}\Gamma_{N}$ (4.37)
where $c_{m}=\mathcal{O}(1)$ is a numerical coefficient. Combining all of these with Eq. (3.82) for $\Gamma_{N_{1}}/H_{\rm eq}$, the present value of $Y_{B,\rm res}$ in Eq. (4.36) is given by
$Y_{B,\rm res}$ $\sim$ $\frac{\pi^{2}{\tilde{\delta}}_{23}c_{m}^{2}}{700\left(c_{m}^{2}+1\right)\lambda_{\phi}^{1/4}{\rm B}_{1}\sqrt{{\rm B}_{N}}}\left(\frac{g_{*}(T_{\gamma,N})}{15\sqrt{2}}\right)^{1/2}\frac{\sqrt{m_{\nu_{1}}\phi_{0}}v_{h}}{m_{\nu}M_{\rm P}}$ (4.38)
$\simeq$ $\frac{3.0\times 10^{-6}\times{\tilde{\delta}}_{23}c_{m}^{2}}{\left(c_{m}^{2}+1\right){\rm B}_{1}\sqrt{{\rm B}_{N}\xi_{\phi}}}\left(\frac{10^{3}m_{\nu_{1}}}{m_{\nu}}\right)^{1/2}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{1/2}\ .$
Therefore, as shown in Fig. 13, even for ${\tilde{\delta}}_{23}\ll 1$ we can obtain a right amount of baryon asymmetry consistent with observations in the present universe for $\phi_{0}\sim\phi_{0}^{\rm ref}$ which saturates the relic density of dark matter consisting of axi-majorons.
Figure 13: The case of resonant leptogenesis for the late time baryon number asymmetry, $Y_{B,\rm res}$ in Eq. (4.38) matching observations as a function of $m_{\nu_{1}}$ and ${\rm B_{1}}$ for ${\tilde{\delta}}_{23}=10^{-5}$(solid red), $5\times 10^{-5}$(dashed red) wtih $c_{m}=1,\ b_{N}=0.15$, $\xi_{\phi}=50$, $\phi_{0}=\phi_{0}^{\rm ref}$, and $m_{\nu_{1}}=10^{-3}m_{\nu}$. The scheme of the blue region and the blue lines is the same as the one in Fig. 9.
5 Dark matter and dark radiation
In our scenario, two types of axi-majorons appear: hot and cold axi-majorons. Hot axi-majorons are produced in decays of the Peccei-Quinn inflaton as discussed in Section 3. At the epoch of the late time decay of the lightest RHN mass-eigenstate, the energy density of axi-majorons to the radiation density of SM particles is given by
$\frac{\rho_{a}}{\rho_{r}}$ $=$ $\left(\frac{H_{\rm d}}{H_{1,\rm eq}}\right)^{2/3}$ (5.1)
with $H_{\rm d}=(2/3)\Gamma_{N_{1}}$. Depending $\Gamma_{N_{1}}\ll H_{1,\rm eq}$ which determines the duration of an early matter-domination era, the contribution of hot axi-majorons to the total energy density of the universe can be still sizable and contributes to the effective number of extra neutrino species, $\Delta N_{\rm eff}$, as a component of dark radiation. In the present universe, it is expected to be
$\Delta N_{\rm eff}$ $\simeq$ $0.48\times\left(\frac{100}{g_{*}(T_{\gamma,\rm d})}\right)^{1/3}\left(\frac{10\Gamma_{N_{1}}}{H_{1,\rm eq}}\right)^{2/3}\ ,$ (5.2)
where $T_{\gamma,\rm d}$ is the photon temperatures at the epoch of $N_{1}$’s decay. It may help to ameliorate the tension in observations for the expansion rate of the present universe [30].
Cold axi-majorons in our scenario are produced only via the misalignment mechanism, since $U(1)_{\rm PQ}$ is already broken during inflation and never restored. It is the only component of cold dark matter in our scenario. In the minimal DFSZ axion scenario we are considering here, the axion decay constant is given by $f_{a}=\phi_{0}/c_{a}$ with $c_{a}=6$ being the QCD-anomaly coefficient of $U(1)_{\rm PQ}$ symmetry with the charge of PQ-field $\Phi$ being normalized to be unity. In conventional scenarios where the initial kinetic energy of the axion is ignored, for $\theta_{\rm mis}=\mathcal{O}(1)$ as the misalignment angle, a right amount of dark matter is obtained for $f_{a}\approx 3\times 10^{11}{\rm GeV}$ [66]. If axi-majoron has a sufficiently large kinetic energy to delay the onset of its low energy coherent oscillation due to angular misalignment, there would be a chance of the kinetic misalignment [67] even for $\phi_{0}$ much smaller than $\phi_{0}^{\rm ref}$. It requires $Y_{\Phi}=\mathcal{O}(10)\times\left(\phi_{0}/10^{9}{\rm GeV}\right)$ as the ratio of the PQ-number density to the entropy density of the SM particles around the epoch of axi-majoron oscillations in conventional scenarios. However, in our scenario, as shown in Eq. (4.5), the angular motion is limited as $Y_{\Phi,e}\lesssim\mathcal{O}(1)$ for $\alpha\lesssim\mathcal{O}(10^{-2})$ which is necessary to obtain sufficiently large motion of axi-majoron after inflation (see Eq. (4.5)). Hence, if axi-majorons are to be the main component of dark matter, the symmetry-breaking scale is constrained as $\phi_{0}\sim\phi_{0}^{\rm ref}$ for $\mathcal{O}(1)$ misalignment angle.
The amount of dark radiation is determined by the duration of the matter-domination (MD) era driven by the lightest RHN, as shown in Eq. (5.2). The presence of an early MD-era reduces the $e$-folds associated with cosmic scales relative to the case of the standard cosmology. As a result, the spectral index in our scenario is expected to be smaller than in the standard case. The difference relative to the expectation in the standard cosmology is small, but this implies that, if some amount of dark radiation is detected in future experiments, in principle, it can be used for a consistency check of our scenario with the spectral index of density perturbations, although it will strongly depend on the sensitivity of the experiments. Also, such an early MD-era causes dilution of pre-existing particles due to the entropy release at the end of the era. Specially, the nearly scale-invariant inflationary gravitational waves would be diluted for cosmic scales which were inside the horizon at the epoch of the decay of the dominating particle. If detected, firstly, the inflationary gravitational waves provide information of the energy scale during inflation, i.e., $H_{I}$ in Eq. (3.20). When combined with inflationary observables, $\lambda_{\phi}$ and $\xi_{\phi}$ can be nearly fixed, modulo the minor variation due to the possible angular motion during inflation when $\alpha$ is somewhat sizable. Additionally, the spectra of the primordial gravitational waves at the present universe will have spectral changes at two frequencies: If the lightest RHNs are responsible for recovering the standard thermal background, the characteristic frequencies are given by
$f_{{\rm eq},0}$ $=$ $f_{N_{1},0}\left(\frac{H_{\rm eq}}{H_{\rm d}}\right)^{2/3},$ (5.3)
$f_{N_{1},0}$ $=$ $f_{N_{1},\rm d}\left(\frac{g_{*S}(T_{\gamma,0})}{g_{*S}(T_{\gamma,\rm d})}\right)^{1/3}\frac{T_{\gamma,0}}{T_{\gamma,\rm d}}$ (5.4)
$\simeq$ $2.7\times 10^{-6}{\rm Hz}\times\left(\frac{H_{\rm d}}{H_{\rm ew}}\right)^{1/2}\ .$
In the second line of Eq. (5.4), we used $f_{N_{1},\rm d}=H_{\rm d}/2\pi$ with $H_{\rm d}=(2/3)\Gamma_{N_{1}}$, and $H_{\rm ew}$ is the expansion rate of the universe at the epoch of the EWPT with photon temperature $T_{\gamma,\rm ew}=100{\rm GeV}$. These frequencies are associated with the beginning (Eq. (5.3)) and the end(Eq. (5.4)) of the matter-domination era. All modes of $f>f_{{\rm eq},0}$ are suppressed by the factor $\left(H_{\rm d}/H_{\rm eq}\right)^{2/3}$ while modes of $f<f_{N_{1},0}$ are not affected. Once such a feature of GWs is detected in some future experiments sensitive to the energy scale of inflation in our scenario(corresponding to $\Omega_{\rm GW}h^{2}\sim\mathcal{O}(10^{-17})$ as the fractional energy of GWs at the present universe), it would provide information of $H_{\rm eq}$ and $\Gamma_{N_{1}}$ separately [68], as long as $\Gamma_{N_{1}}/H_{\rm eq}$ is not much smaller than unity. Even in the case only $f_{N_{1},0}$ is detected, if available, $\Delta N_{\rm eff}$ can be used to determine $H_{\rm eq}$. Then, since dark matter relic density would nearly fix $\phi_{0}$, it may be possible to extract information of $m_{\nu_{1}}$ and $m_{N_{1}}$. To our knowledge, this is a unique way of probing those mass parameters in the seesaw sector.
6 Conclusions
In this work, we showed the details of ‘the minimal cosmological standard model (MCSM)’ which we proposed recently in Ref. [1]. The model is a minimal scenario beyond the standard model, which includes the seesaw sector too in the form of the singlet majoron scenario. Scale-invariance was imposed for both of the matter and the gravity sectors. the Planck scale emerges dynamically from the spontaneous breaking of scale-invariance. The global Peccei-Quinn symmetry, $U(1)_{\rm PQ}$, was imposed on the matter sector, but it was assumed to be broken in the gravity sector in a scale-invariant manner. The model is a combination of DFSZ axion model and the singlet majoron scenario, in which Peccei-Quinn field not only provides the axion solution but also is responsible for generation of the masses of right-handed neutrinos for the seesaw mechanism.
We showed that our scenario provides a simple unified framework simultaneously addressing the following phenomenological/cosmological/theoretical puzzles: (i) the origin of scales, (ii) a period of primordial inflation, (iii) baryogenesis, (iv) tiny neutrino masses, (v) dark matter, and (vi) strong CP-problem. Primordial inflation in our scenario is realized along the Peccei-Qinn field direction thanks to a non-minimal gravitational interaction and scale-invariance. Thanks to the symmetry-breaking non-minimal coupling, in general angular motion of the inflaton arises, and can generate a large PQ-number asymmetry at the end of inflation. It is nothing but the way of generating particle asymmetry in Affleck-Dine mechanism of baryogenesis, which is naturally realized in our setup. Depending on parameters, the asymmetry can be transferred to the visible sector through right-handed neutrinos in various ways. It can be through a non-thermal decay of the lightest RHNs which have a helicity-asymmetry induced by the asymmetry of PQ-number. This is a kind of usual Affleck-Dine scenario, although details are somewhat different in our scenario. Alternatively, the asymmetry in the visible sector can be obtained via thermal processes, as in the case of spontaneous baryogenesis mechanism. In this case, the remnant angular motion of axi-majoron after the decay of the Peccei-Quinn inflaton provides the effective chemical potential to the fields charged under $U(1)_{\rm PQ}$-symmetry. The third possibility is the resonant leptogenesis from the decay of nearly degenerate heavy RHN mass-eigenstates. This case is relevant when the initial PQ-number asymmetry after inflation is not enough for neither AD-mechanism nor spontaneous baryogenesis.
Cold dark matter is purely axi-majorons produced only through the misalignment mechanism unless the mass of the lightest left-handed neutrino is essentially zero. Kinetic misalignment mechanism does not work, since the kick of the axi-majoron at the end of inflation is limited by axi-majoron’s initial position and velocity during inflation. Hence, the symmetry-breaking scale of $U(1)_{\rm PQ}$, i.e., $\phi_{0}$ is constrained to be of $\mathcal{O}(10^{12}){\rm GeV}$. A sizable amount of dark radiation can be obtained naturally, and may be helpful for ameliorating the tension between observations for the expansion rate of the present universe.
In regard to axion solution, the symmetry-breaking term in the gravity sector does not invalidate the solution, since its contribution to the axion potential is proportional to the total potential energy of the universe, and becomes negligible in the late universe. Although our scenario is a DSFZ axion model having more than one domain, it does not have a domain wall problem thanks to the sizable non-zero angular motion and weak interactions of Peccei-Quinn field to RHNs. Iso-curvature perturbations of the axi-majoron are also suppressed because of the following reasons: (i) the magnitude of Peccei-Quinn field during inflation is comparable to or larger than the Planck scale, and (ii) the curvature along the angular direction is either positive or much smaller than the expansion rate when the relevant cosmological scales exit the horizon. The axion-quality problem may also be absent in the presence of Gauss-Bonnet term.
Our scenario is a very simple minimal extension of the SM, but it provides a unified picture of the unknown history of the universe from inflation to the BBN epoch, thanks to a scale-invariant $U(1)_{\rm PQ}$-breaking non-minimal gravitational interaction. This model predicts a presence of a short early matter-domination era. It determines the amount of axi-majoron dark radiation, and leaves a characteristic fingerprint on inflationary primordial gravitational waves. If such a feature is detected in future experiments, and once combined with inflationary observables, the amount of dark radiation and the spectra of gravitational waves can provide information of some mass parameters of the left-handed and the right-handed neutrinos. Although complicated, it may end up being the unique way of obtaining those physical quantities.
Acknowledgments
This work was supported by the National Research Foundation of Korea grants by the Korea government: 2017R1D1A1B06035959 (W.I.P.), 2022R1A4A5030362 (W.I.P.), NRF-2019R1A2C3005009 (P.K.) and KIAS Individual Grants under Grant No. PG021403 (P.K.). It was also supported by the Spanish grants CIPROM/2021/054 (Generalitat Valenciana), PID2020-113775GB-I00 (AEI/10.13039/501100011033) (G.B. and W.I.P.) and by the European ITN project HIDDeN (H2020-MSCA-ITN-2019/860881-HIDDeN) (G.B.).
Appendix A Metric component in a various basis
For a fixed Planck scale with the SM Higgs field ignored, we define the frame function as
$\Omega\equiv 1+\frac{\xi_{\phi}\phi^{2}}{M_{\rm P}^{2}}\left[1+\alpha\cos\left(2\theta_{\Phi}\right)\right]=1+\frac{\xi_{r}\phi_{r}^{2}}{M_{\rm P}^{2}}+\frac{\xi_{i}\phi_{i}^{2}}{M_{\rm P}^{2}}\ ,$ (A.1)
where $\alpha\equiv\xi_{a}/\xi_{\phi}$, and the condition
$\xi_{r}>\xi_{i}\quad\Leftrightarrow\quad\xi_{\phi}>\xi_{a}$ (A.2)
is required for $\Omega$ to be positive definite even at the large field region of $\xi_{\phi}^{1/2}\phi>M_{\rm P}$ which is of our interest. The symmetric field-space metric $\mathcal{K}_{ab}$ is given by
$\mathcal{K}_{ab}=\Omega^{-2}\left[\Omega\delta_{ab}+\frac{3M_{\rm P}^{2}}{2}\frac{\partial\Omega}{\partial\phi_{a}}\frac{\partial\Omega}{\partial\phi_{b}}\right]\equiv\Omega^{-2}K_{ab}\ .$ (A.3)
In the basis of ($\phi_{r},\phi_{i}$),
$\frac{\partial\Omega}{\partial\phi_{r}}$ $=$ $\frac{2\xi_{r}}{M_{\rm P}^{2}}\phi_{r}\ ,$ (A.4)
$\frac{\partial\Omega}{\partial\phi_{i}}$ $=$ $\frac{2\xi_{i}}{M_{\rm P}^{2}}\phi_{i}\ ,$ (A.5)
and one finds
$K_{11}$ $=$ $1+\frac{\xi_{r}\left(1+6\xi_{r}\right)\phi_{r}^{2}}{M_{\rm P}^{2}}+\frac{\xi_{i}\phi_{i}^{2}}{M_{\rm P^{2}}}\ ,$ (A.6)
$K_{22}$ $=$ $1+\frac{\xi_{r}\phi_{r}^{2}}{M_{\rm P}^{2}}+\frac{\xi_{i}\left(1+6\xi_{i}\right)\phi_{i}^{2}}{M_{\rm P^{2}}}\ ,$ (A.7)
$K_{12}$ $=$ $\frac{6\xi_{r}\xi_{i}\phi_{r}\phi_{i}}{M_{\rm P}^{2}}\ .$ (A.8)
In the ($\phi,\theta$)-basis,
$\frac{\partial\Omega}{\partial\phi}$ $=$ $\frac{2\xi_{\phi}}{M_{\rm P}^{2}}\phi\left[1+\alpha\cos\left(2\theta_{\Phi}\right)\right]\ ,$ (A.9)
$\frac{\partial\Omega}{\phi\partial\theta}$ $=$ $-\frac{2\xi_{\phi}}{M_{\rm P}^{2}}\phi\alpha\sin\left(2\theta_{\Phi}\right)\ ,$ (A.10)
and
$K_{11}$ $=$ $\Omega+6\left(\frac{\xi_{\phi}\phi}{M_{\rm P}}\right)^{2}\left[1+\alpha\cos\left(2\theta_{\Phi}\right)\right]^{2}\ ,$ (A.11)
$K_{22}$ $=$ $\Omega+6\left(\frac{\xi_{\phi}\phi}{M_{\rm P}}\right)^{2}\left[\alpha\sin\left(2\theta_{\Phi}\right)\right]^{2}\ ,$ (A.12)
$K_{12}$ $=$ $-6\left(\frac{\xi_{\phi}\phi}{M_{\rm P}}\right)^{2}\left[1+\alpha\cos\left(2\theta_{\Phi}\right)\right]\left[\alpha\sin\left(2\theta_{\Phi}\right)\right]\ .$ (A.13)
Appendix B Derivatives of the Einstein frame potential w.r.t. the canonical fields
In the large field region of $\xi_{\phi}^{1/2}\phi\gg M_{\rm P}$, in terms of the approximate canonical fields, the potential is read as
$U(\varphi_{1},\varphi_{2})$ $=$ $\frac{\lambda_{\Phi}M_{\rm P}^{4}}{4\xi_{\phi}^{2}}\left[1+\alpha\cos\left(\frac{2\sqrt{\xi_{\phi}}\varphi_{2}}{M_{\rm P}}\right)+\exp\left[-2\sqrt{\frac{\xi_{\phi}}{1+6\xi_{\phi}}}\frac{\varphi_{1}}{M_{\rm P}}\right]\right]^{-2}\ .$ (B.1)
Let us define $A_{1}\equiv 2\sqrt{\xi_{\phi}/(1+6\xi_{\phi})}$ and $A_{2}\equiv 2\sqrt{\xi_{\phi}}$ for notational convenience. Then, derivatives of the potential are found to be
$\frac{\partial U}{\partial\varphi_{1}}$ $=$ $\frac{2A_{1}U/M_{\rm P}}{\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]}$ (B.2)
$\frac{\partial^{2}U}{\partial\varphi_{1}^{2}}$ $=$ $\frac{2A_{1}^{2}\left[2-\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]U/M_{\rm P}^{2}}{\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{2}}$ (B.3)
$\frac{\partial^{3}U}{\partial\varphi_{1}^{3}}$ $=$ $\frac{2A_{1}^{3}\left[4-7\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)+\exp\left[\frac{2A_{1}\varphi_{1}}{M_{\rm P}}\right]\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)^{2}\right]}{\left(U/M_{\rm P}^{3}\right)^{-1}\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{3}}$ (B.4)
$\frac{\partial U}{\partial\varphi_{2}}$ $=$ $\frac{2A_{2}\alpha\sin\left(A_{2}\varphi_{2}/M_{\rm P}\right)U/M_{\rm P}}{\left[1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)+\exp\left[-\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]}$ (B.5)
$\frac{\partial^{2}U}{\partial\varphi_{2}^{2}}$ $=$ $\frac{2A_{2}^{2}\alpha\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\left[\left(1+\left(1-\alpha\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right)\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)+2\alpha\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]}{\left(U/M_{\rm P}^{2}\right)^{-1}\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{2}}$ (B.6)
$\frac{\partial^{3}U}{\partial\varphi_{2}^{3}}$ $=$ $\frac{2A_{2}^{3}\alpha\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]f(\varphi_{1},\varphi_{2})\sin\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)}{\left(U/M_{\rm P}^{3}\right)^{-1}\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{3}}$ (B.7)
$\frac{\partial^{2}U}{\partial\varphi_{1}\partial\varphi_{2}}$ $=$ $\frac{6A_{1}A_{2}\alpha\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\sin\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)U/M_{\rm P}^{2}}{\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{2}}\ ,$ (B.8)
where
$f(\varphi_{1},\varphi_{2})$$\equiv$ $-1+\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\left[-2+7\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right]$ (B.9)
$+\exp\left[\frac{2A_{1}\varphi_{1}}{M_{\rm P}}\right]\left[-1+10\alpha^{2}+7\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)-2\alpha^{2}\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right]\ .$
For $\alpha\lll 1$ they are approximated as
$\frac{\partial U}{\partial\varphi_{1}}$ $=$ $2A_{1}\exp\left[-\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\frac{U}{M_{\rm P}}$ (B.10)
$\frac{\partial^{2}U}{\partial\varphi_{1}^{2}}$ $=$ $-2A_{1}^{2}\exp\left[-\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\frac{U}{M_{\rm P}^{2}}$ (B.11)
$\frac{\partial^{3}U}{\partial\varphi_{1}^{3}}$ $=$ $2A_{1}^{3}\left(-7+\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right)\exp\left[-2\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\frac{U}{M_{\rm P}^{3}}$ (B.12)
$\frac{\partial U}{\partial\varphi_{2}}$ $=$ $2A_{2}\alpha\sin\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\frac{U}{M_{\rm P}}$ (B.13)
$\frac{\partial^{2}U}{\partial\varphi_{2}^{2}}$ $=$ $2A_{2}^{2}\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\frac{U}{M_{\rm P}^{2}}$ (B.14)
$\frac{\partial^{3}U}{\partial\varphi_{2}^{3}}$ $=$ $-2A_{2}^{3}\alpha\sin\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\frac{U}{M_{\rm P}^{3}}\xrightarrow[]{\theta_{\Phi}\ll 1}-4\alpha\theta_{\Phi}\frac{\lambda_{\Phi}M_{\rm P}}{\sqrt{\xi_{\phi}}}=-8\sqrt{3/\lambda_{\Phi}}\alpha\theta_{\Phi}H_{I}\ .$ (B.15)
Appendix C Decay rate of right-handed neutrinos
For reader’s convenience, here we collect the partial decay rates of RHNs to 2-body final states from Ref. [69, 70].
  • $m_{N}>m_{W}$: The decay rate is given by [69]
    $\Gamma^{N}=\frac{y_{\nu}^{2}m_{N}}{64\pi}\times\sum_{i}n_{i}\Theta(m_{N}-m_{i})\left(1-\frac{m_{i}^{2}}{m_{N}^{2}}\right)^{2}\left(1+\frac{2m_{i}^{2}}{m_{N}^{2}}\right)^{n_{V,i}}$ (C.1)
    where $i=(h,Z,W)$, $n_{i}=(1,1,2)$ is the number of channels, $n_{V,i}=(0,1,1)$ is the vector boson number, and $\Theta$ is the Heaviside step function.
  • $m_{N}<m_{W}$: The total 2-body rate of $N$ is given by [70]
    $\Gamma^{N}=\frac{|U_{\alpha}|^{2}}{32\pi}G_{F}^{2}m_{N,0}^{3}\times\sum_{i}\mathcal{F}_{i}$ (C.2)
    where
    $|U_{\alpha}|^{2}=\frac{y_{\nu}^{2}v_{h}^{2}}{4m_{N}^{2}}=\frac{m_{\nu}}{2m_{N}}$ (C.3)
    and $\mathcal{F}_{i}$ for each decay channel is given by
    $\pi^{0}+\nu_{\alpha}$ ::: $f_{\pi}^{2}\left(1-\frac{m_{\pi}^{2}}{m_{N}^{2}}\right)^{2}$ (C.4)
    $H^{+}\ell_{\alpha}^{-}$ ::: $2f_{H}^{2}|V_{H}|^{2}\left[\left(1-\frac{m_{\ell}^{2}}{m_{N}^{2}}\right)^{2}-\frac{m_{H}^{2}}{m_{N}^{2}}\left(1+\frac{m_{\ell}^{2}}{m_{N}^{2}}\right)\right]$ (C.5)
    ×$\times\sqrt{\left(1-\frac{(m_{H}-m_{\ell})^{2}}{m_{N}^{2}}\right)\left(1-\frac{(m_{H}+m_{\ell})^{2}}{m_{N}^{2}}\right)}$
    $\eta+\nu_{\alpha}$ ::: $f_{\eta}^{2}\left(1-\frac{m_{\eta}^{2}}{m_{N}^{2}}\right)^{2}$ (C.6)
    $\eta^{\prime}+\nu_{\alpha}$ ::: $f_{\eta^{\prime}}^{2}\left(1-\frac{m_{\eta^{\prime}}^{2}}{m_{N}^{2}}\right)^{2}$ (C.7)
    $\rho^{+}\ell_{\alpha}^{-}$ ::: $4\frac{g_{\rho}^{2}}{m_{\rho}^{2}}|V_{ud}|^{2}\left[\left(1-\frac{m_{\ell}^{2}}{m_{N}^{2}}\right)^{2}-\frac{m_{\rho}^{2}}{m_{N}^{2}}\left(1+\frac{m_{\ell}^{2}-2m_{\rho}^{2}}{m_{N}^{2}}\right)\right]$ (C.8)
    ×$\times\sqrt{\left(1-\frac{(m_{\rho}-m_{\ell})^{2}}{m_{N}^{2}}\right)\left(1-\frac{(m_{\rho}+m_{\ell})^{2}}{m_{N}^{2}}\right)}$
    $\rho^{0}+\nu_{\alpha}$ ::: $2\frac{g_{\rho}^{2}}{m_{\rho}^{2}}\left(1-\frac{m_{\rho}^{2}}{m_{N}^{2}}\right)^{2}\left(1+\frac{2m_{\rho}^{2}}{m_{N}^{2}}\right)$ (C.9)
    with
    $f_{\pi}=130{\rm MeV},f_{\eta}=1.2f_{\pi},\ f_{\eta^{\prime}}=-0.45f_{\pi},\ g_{\rho}=0.102{\rm GeV}^{2},$ (C.10)
    and $f_{H}$ and $V_{H}$ are the decay constants and the CKM matrix associated with the final state mesons, $H^{+}=\pi^{+},K^{+},D^{+},D_{s}^{+},B^{+},B_{c}^{+}$ [70].
Fig 14 shows the total decay rate, $\Gamma_{N_{i}}$, of the mass-eigenstate $N_{i}$ as a function of the mass, $m_{N_{i}}$, where $y_{\nu}$ was assumed to have the seesaw relation in Eq. (3.57).
Figure 14: Decay rate of a RHN mass-eigenstate as a function of its mass $m_{N}$. The vertical dotted line corresponds to $m_{N}=m_{W}$. The horizontal line is a reference line of the expansion rate at the electroweak scale background temperature in the standard radiation-domination universe. It is assumed that RHN decay is flavor-blind for left-handed neutrinos and charged leptons separately. Hence, for a neutrino channel and lepton channel a factor three was multiplied, respectively.